### v1

# Adaptive Gradient Calibration for Single-Positive Multi-Label Learning in Remote Sensing Image Scene Classification

Abstract
Multi-label classification (MLC) offers a more comprehensive semantic understanding of Remote Sensing (RS) imagery compared to traditional single-label classification (SLC). However, obtaining complete annotations for MLC is particularly challenging due to the complexity and high cost of the labeling process. As a practical alternative, single-positive multi-label learning (SPML) has emerged, where each image is annotated with only one relevant label, and the model is expected to recover the full set of labels. While scalable, SPML introduces significant supervision ambiguity, demanding specialized solutions for model training. Although various SPML methods have been proposed in the computer vision domain, research in the RS context remains limited. To bridge this gap, we propose Adaptive Gradient Calibration (AdaGC), a novel and generalizable SPML framework tailored to RS imagery. AdaGC adopts a gradient calibration (GC) mechanism combined with Mixup and a dual exponential moving average (EMA) module for robust pseudo-label generation. To maximize AdaGC’s effectiveness, we introduce a simple yet theoretically grounded indicator to adaptively trigger GC after an initial warm-up stage based on training dynamics, thereby guaranteeing the effectiveness of GC in mitigating overfitting to label noise. Extensive experiments on two benchmark RS datasets under two distinct label noise types demonstrate that AdaGC achieves state-of-the-art (SOTA) performance while maintaining strong robustness across diverse settings.
I Introduction
With the rapid development of satellite missions, vast amounts of remote sensing (RS) imagery have become increasingly accessible, serving as a critical data source for numerous applications [1, 2]. As a fundamental task, image classification plays a vital role in connecting imagery with downstream applications in the RS domain [3, 4, 5]. However, most existing RS studies adopt the single-label classification (SLC) paradigm, assigning a single semantic label to each image [6, 7], which oversimplifies the inherent complexity of real-world RS scenes [8, 9]. In practice, RS images usually contain diverse land cover types, objects, or functional zones within a single scene. As illustrated in Fig. 1, very-high resolution (VHR) imagery often captures land use and land cover information as well as fine-grained objects such as buildings, vehicles, and vegetation within a limited area. Lower-resolution imagery often spans heterogeneous regions, such as agricultural mosaics or urban–rural transitions. In both cases, multi-label classification (MLC) offers a more informative, comprehensive, and realistic semantic representation of RS imagery compared to traditional SLC methods.
However, MLC faces a major challenge in acquiring complete and accurate annotations due to the difficulty and high cost in the annotation process [10]. This challenge is particularly critical in the deep learning (DL) era, where large-scale labelled datasets are essential [11]. A practical compromise is to annotate each image with only one relevant label, even when multiple classes are present, which leads to the single-positive multi-label learning (SPML) setting [12]. In this setting, the model is weakly supervised with one positive label per image, yet it is expected to recover the complete set of relevant labels. Supervision in SPML is inherently weak, characterized by missing labels. In practice, even expert annotators often miss relevant labels in multi-label settings, making the false negative issue pervasive in MLC. SPML can thus be viewed as an extreme case of this problem and also offers a promising solution to completing label sets in single-label annotation scenarios.
Refer to caption Refer to caption Refer to caption
(a) Single-label: court Multi labels: buildings, cars, pavement, court, grass, trees (b) Single-label: arable land (dominant) Multi-labels: arable land, urban fabric, inland waters, complex cultivation patterns
Figure 1: Single- and multi-label annotation examples from (a) AID-multilabel [13] and (b) refined BigEarthNet [14] datasets, where the corresponding CLC mask is presented for reference. Compared to the single-class labels, the multi-label annotations can more comprehensively describe the scene.
Although SPML is promising in scalability and annotation efficiency, it also introduces significant ambiguity that can bias the learning process. A common baseline approach is the AN strategy [12], which treats all unobserved labels as negatives, unavoidably introducing bias into learning. In the computer vision (CV) domain, several strategies have been proposed to address its limitations, such as noise-robust regularization terms, online label-correction/pseudo-labeling methods, and interpretability-driven frameworks [12, 15, 16, 17, 18, 19, 20]. Despite the advancements in CV, the SPML problem remains underexplored in RS. SPML is particularly valuable for RS image classification, where annotation typically requires greater domain expertise and incurs higher costs compared to natural image datasets. These challenges make the application of SPML in RS more necessary and also more difficult.
When directly applying existing SPML methods to RS data, several challenges arise. One major issue is the reliability of pseudo-labeling methods. VHR RS images often include small, sparsely distributed targets, while medium-to-low-resolution imagery tends to capture broader regional semantics. Both cases can easily result in low inter-class separability and high intra-class variability, hindering the effectiveness of simple pseudo-labeling strategies such as confidence thresholding. Furthermore, the distinct characteristics of MLC versus multi-class classification (MCC) suggest that commonly used learning-from-noisy-labels strategies, which are often designed for MCC, should be carefully tailored to SPML. Moreover, although early learning theory indicates that models gradually reach a peak generalization performance in the early stage of training before overfitting to label noise [21, 22], most existing SPML methods rely on predefined warm-up schedules and overlook this dynamic behaviour. This, to some extent, limits the model’s capability in pseudo-labeling.
In this work, we propose a novel and generalizable framework for SPML in RS, termed Adaptive Gradient Calibration (AdaGC), to address the aforementioned challenges. The framework consists of three key components. First, we introduce a dual exponential moving average (EMA) module for robust pseudo-label estimation, which leverages the consensus between the temporally smoothed predictions of the student model via EMA and the outputs of the teacher model, whose weights are updated via EMA of the student’s parameters. Second, the generated pseudo-labels are used to guide a Gradient Calibration (GC) mechanism inspired by Early Learning Regularization (ELR) [21], which is originally designed for MCC, to alleviate overfitting to false negatives. To further improve generalizability, we incorporate the effective data augmentation technique Mixup [15] within the GC process. Finally, to ensure the effectiveness of this calibration process, we conduct an in-depth analysis of learning dynamics in SPML and propose a simple yet theoretically grounded indicator for adaptive activation of GC after an initial warm-up stage. This mechanism is crucial for avoiding the negative influence of low-quality pseudo-labels on model optimization. We evaluate the AdaGC on two benchmark RS multi-label datasets with two single-positive label simulation ways. Extensive experiments demonstrate the effectiveness and robustness of AdaGC.
The main contributions of this paper are summarized as follows:
  • We propose a two-stage SPML framework for RS imagery based on a student–teacher architecture, incorporating an early learning detection strategy that guides the activation of GC to combat overfitting to label noise. Furthermore, we provide theoretical analyses that support the effectiveness of our approach in SPML scenarios.
  • We introduce a dual-EMA pseudo-label generation strategy that enhances label completeness and reliability by leveraging temporal prediction fusion.
  • We perform extensive empirical evaluations on two benchmark RS datasets, covering high- and low-resolution imagery, and two single positive label noise simulation strategies (Random and Dominant). We also provide a comprehensive benchmarking of existing SPML methods, offering valuable insights for future research in this domain.
The remainder of this paper is organized as follows. Section II reviews existing works on MLC, with a focus on noisy labels and SPML. Section III details the proposed method along with the theoretical analyses. The empirical evaluation is presented in Section IV, with the conclusion in Section V.
II Related Works
This section reviews advancements in MLC, first focusing on its application in RS and then discussing in general the challenges posed by noisy labels. Finally, we delve into SPML, detailing various methods developed in CV to address its unique difficulties.
II-A Remote Sensing Image Multi-label Classification
Early DL research in MLC for RS imagery primarily focused on high-resolution data and employed Convolutional Neural Networks for feature extraction, followed by classifiers such as Radial Basis Function Neural Networks [23] or Structured Support Vector Machines [24]. These approaches often relied on conventional transfer learning, using CNNs pre-trained on generic CV datasets (e.g., ImageNet [25]) as fixed feature extractors. However, the significant differences between natural CV scenes and complex RS imagery can lead to suboptimal classification performance, limiting the effectiveness of such direct transfer.
To mitigate the limitations of pre-trained models, alternative strategies have been proposed. For instance, Stivaktakis et al. [26] introduced a data augmentation scheme to facilitate end-to-end training of shallow CNNs, along with replacing the softmax layer with a sigmoid function to accommodate multi-label outputs. While this direct adaptation remains a common practice in multi-label scenarios, it may result in the imprecise identification of multiple classes. More sophisticated approaches integrate sequential neural network architectures with CNNs to enhance multi-label scene classification. Hua et al. [8] proposed a class-wise attention-based Recurrent Neural Network (RNN) to explicitly model the co-occurrence relationships among multiple classes, thus generating class predictions sequentially. Similarly, an attention-aware label relational reasoning network was introduced in [13] to localize discriminative regions and characterize inter-label relationships based on feature maps. Alshehri et al. [27] presented an encoder-decoder neural network architecture, where the encoder includes a squeeze-and-excitation layer to model channel-wise dependencies, and the RNN-based decoder serves as an adaptive spatial attention module. These attention-driven strategies have demonstrated considerable success in identifying informative areas through attention maps derived from convolutional features.
II-B Multi-label Classification with Noisy Labels
A major hurdle in MLC in RS is the difficulty in obtaining exhaustive multi-label annotations, which often results in noisy labels. This noise can be categorized as either missing labels (subtractive noise), where a class present in the considered image is not annotated, or as wrong labels (additive noise), where an absent class is incorrectly annotated [10]. Such multi-label noise, which can be introduced through zero-cost annotation (e.g., using thematic products), can severely impair model training. Consequently, developing methods robust to multi-label noise in RS is an active research area, usually categorized as weakly supervised multi-label learning (WSML) [28].
Methods for handling noisy labels in multi-label settings often draw inspiration from related fields, such as positive-unlabeled learning (PUL) [29, 30, 31] and noisy multi-class learning [32]. Common strategies involve developing robust loss functions that are less sensitive to label noise or class imbalance [33]. Another approach is pseudo-labeling, where likely labels for unobserved or unknown annotations are inferred to provide additional training signals [34]. The “memorization effect,” where models tend to learn clean labels early but later start to memorize noisy ones, has been widely observed in noisy MCC. ELR [21] was developed to specifically counteract this, by adjusting the influence of gradients from potentially noisy samples, particularly in the early stages of training. Burgert et al. [10] later extended it to MLC in a straightforward manner and empirically demonstrated its effectiveness. Additional strategies include sample selection, which aims to identify and potentially exclude samples with noisy labels [35], and label correction/refurbishment, which attempts to correct the noisy labels (e.g., using collaborative learning frameworks [36, 37]).
Other methods, such as those in the large loss (LL) [16] family, including LL-R, LL-Ct, and LL-Cp, are developed to recognize and handle mislabeled samples with large losses, based on the observation that noisy labels often lead to high loss values. Specifically, LL-R rejects samples with large losses, LL-Ct applies temporary corrections, and LL-Cp employs permanent corrections. BoostLU [19] complements LL by enhancing predicted positive regions via a function applied to class activation maps, further improving the robustness.
In WSML for RS, subtractive noise, i.e., missing true labels, is a prevalent issue. This often occurs because annotators, especially when facing complex imagery or large datasets, opt to label only the dominant class(es) rather than exhaustively identify all present categories. This poses a significant challenge because an unannotated class doesn’t necessarily mean its absence, leading to the extreme case of subtractive noise known as SPML. This scenario is both challenging and increasingly common in practical RS applications, despite being underexplored in the RS literature.
II-C Single Positive Multi-Label Learning
[…]
However, the above-mentioned methods are mainly developed for natural images, and their direct application to RS data often falls short due to the distinct characteristics of RS imagery. Moreover, while pseudo-labeling has shown effective, it typically requires a warm-up stage to ensure label quality. Manually determining this stage can be suboptimal, either due to insufficient learning or overfitting to noisy labels.
III Methodology
Figure 2: Flowchart of the proposed Adaptive Gradient Calibration (AdaGC) method for single positive multi-label learning in remote sensing image classification.
As illustrated in Fig. 2, the proposed method adopts a student–teacher framework, where the teacher model’s weights are updated via EMA. The training pipeline consists of two stages. In the first stage, the student model is trained directly on noisy labels, during which the teacher model’s validation accuracy (with respect to noisy labels) is monitored as an early learning indicator to determine when to activate the second stage— Gradient Calibration (GC). Next, we first define the problem before presenting the technical details of the proposed method.
III-A Problem Definition
In SPML, each image $\mathbf{x}_{i}\in\mathbb{R}^{h\times w\times d}$ is associated with a true label vector $\mathbf{y}_{i}^{*}\in\{0,1\}^{C}$, where multiple labels may be relevant. However, during training, only a single positive label is observed per image. The observed label vector $\mathbf{y}_{i}\in\{0,1\}^{C}$ satisfies the following conditions:
$\mathbf{y}_{i}\preceq\mathbf{y}_{i}^{*},\quad\|\mathbf{y}_{i}\|_{1}=1,$ (1)
where $C$ is the total number of classes and $\preceq$ denotes element-wise no-larger-than comparison. Specifically, for two binary vectors $\mathbf{a},\mathbf{b}\in\{0,1\}^{C}$, $\mathbf{a}\preceq\mathbf{b}$ if $a_{j}\leq b_{j},\forall j$.
The objective is to train a multi-label classifier $f$ parametrized by weights $\theta$ to recover the complete set of labels despite incomplete supervision. The common AN strategy treats all unobserved labels as negatives. This leads to the standard binary cross-entropy (BCE) loss:
$\small\mathcal{L}_{\text{AN}}\left(\theta\right)=-\sum_{i=1}^{N}\sum_{c=1}^{C}\left[y_{i,c}\log p_{i,c}+\left(1-y_{i,c}\right)\log{\left(1-p_{i,c}\right)}\right],$ (2)
where $\mathbf{p}_{i}=\texttt{Sigmoid}(f(\mathbf{x}_{i};\theta))\in[0,1]^{C}$. However, this approach can easily lead to overfitting to label noise by incorrectly penalizing potentially relevant but unobserved labels.
III-B Gradient Calibration
The concept of gradient calibration for handling noisy labels was first proposed in [21] for MCC, where ELR introduces an auxiliary regularization term to constrain the standard cross-entropy loss, thus preventing the memorization of noisy labels. The regularization term for MCC is defined as follows:
$\mathcal{R}_{\text{MCC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\log\left(1-\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle\right),$ (3)
where $n$ is the number of training samples, $\mathbf{p}$ denotes the softmax output vector of the multi-class model, and $\mathbf{t}$ represents the soft pseudo-labels. The gradient of (3) is defined as follows:
$\nabla_{\theta}\mathcal{R}_{\text{MCC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}f\left(\mathbf{x}_{i};\theta\right)\cdot\mathbf{g}_{i},$ (4)
with $\mathbf{g}_{i}\in\mathbb{R}^{C}$, whose values are computed as follows:
$\mathbf{g}_{i,c}=\frac{p_{i,c}}{1-\langle\mathbf{p}_{i},\mathbf{t}_{i}\rangle}\sum_{k=1}^{C}\left(t_{i,k}-t_{i,c}\right)\cdot p_{i,k}.$ (5)
As indicated by (5), the sign of (4) is governed by $t_{i,k}-t_{i,c}$. Specifically, the ELR regularization calibrates the gradients toward negative values when $\mathbf{t}_{i}$ indicates that the class $c$ is the potential correct label. Conversely, if class $c$ is not favored by the pseudo label, the calibration term $g_{i,c}$ remains positive. This behavior is consistent with the gradient direction of the cross-entropy loss, which yields negative gradients for $y_{i,c}=1$ and positive gradients for $y_{i,c}=0$.
Equation 3 holds effective only for multi-class cases with $\|\mathbf{p}_{i}\|_{1}=1$ and $\|\mathbf{t}_{i}\|_{1}=1$. In the multi-label case, directly applying (3) leads to unbounded $\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle$ that also varies significantly across samples due to differing numbers of positive labels. A straightforward adaptation is to extend it to a class-wise binary setting, that is,
$\mathcal{R}_{\text{MLC}}^{binary}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\log\left(1-\left\langle\mathbf{b}_{i,c},\mathbf{t}_{i,c}\right\rangle\right),$ (6)
where $\mathbf{b}_{i,c}=[p_{i,c},1-p_{i,c}]^{\mathsf{T}}$ is the binary prediction vector for sample $i$ at class $c$, and $\mathbf{t}_{i,c}=[t_{i,c},1-t_{i,c}]^{\mathsf{T}}$. In general, negative labels ($y_{i,c}=0$) vastly outnumber positive ones ($y_{i,c}=1$) in multi-label datasets, even though each sample may be associated with multiple classes. Thus, (6) falls ineffective in practice, particularly in SPML settings, since the regularization term is largely dominated by calibration on $y_{i,c}=0$, and overwhelms the contribution from the sparse positive labels.
To address this limitation, we emphasize the gradient calibration on the potential positive labels with the following regularization term:
$\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\log\left(1-p_{i,c}\cdot t_{i,c}\right).$ (7)
We can compute the gradient of (7) as follows:
$\nabla_{\theta}\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\nabla_{\theta}f(\mathbf{x}_{i};\theta)\cdot g_{i,c},$ (8)
where $g_{i,c}$ is defined as follows:
$g_{i,c}=\frac{p_{i,c}-1}{\frac{1}{p_{i,c}\cdot t_{i,c}}-1}\leq 0.$ (9)
We argue that (7) is particularly suitable for the SPML setting, as it specifically penalizes false negatives. Accordingly, we apply GC only to negative labels, excluding its use for positive ones in practice. The final objective is a combination of the binary cross-entropy loss and the GC regularization:
$\mathcal{L}\left(\theta\right)=\mathcal{L}_{\text{AN}}\left(\theta\right)+\lambda\cdot\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right),$ (10)
where $\lambda$ is the weight of the regularization term.
However, applying GC from the beginning may lead the model to predict all classes as positive. Liu et al. [22] show that adaptively triggering correction before overfitting to label noise is crucial for the effectiveness of sample-correction-related techniques. Inspired by this, in Section III-C we introduce a simple early learning detection mechanism to appropriately activate GC. Moreover, as GC performance depends on the quality of the pseudo-labels $\mathbf{t}$, in Section III-D we design a dual-EMA module to ensure the robust generation of pseudo-labels.
III-C Early Learning Detection
Pseudo-labelling methods typically require an initial warm-up phase, whose duration, often determined empirically, critically impacts performance. In SPML, premature pseudo-label initiation introduces additive noise, whereas excessive delay risks overfitting to inherent subtractive label noise, thus limiting false negative detection. Optimal pseudo-labelling activation is crucial for model performance. Monitoring performance metrics on a clean validation set is ideal, as model performance typically improves, plateaus, then declines (indicating overfitting) [40, 41]. The plateau, detectable by the slope of performance metrics, identifies the optimal time to activate pseudo-labelling [22]. However, SPML generally lacks clean validation data, typically offering only single positive labels, akin to the training set. While standard metrics on noisy sets might differ, we demonstrate that the mean average precision (mAP) computed on the noisy validation set mirrors the behavior of mAP* computed on the clean validation set, showing approximately consistent trends throughout training.
The mAP is the macro average of average precision (AP) for each class:
$\text{mAP}=\frac{1}{C}\sum_{c=1}^{C}{\text{AP}_{c}},$ (11)
where $\text{AP}_{c}=\int_{0}^{1}{\text{Prec}_{c}\left(\text{Rec}_{c}\right)d\text{Rec}_{c}}$. For a clean validation set, recall $\text{Rec}^{*}_{c}$ and precision $\text{Prec}^{*}_{c}$ are defined as follows:
$\begin{split}\text{Rec}^{*}_{c}=\frac{\text{TP}_{c}}{\text{P}_{c}}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FN}_{c}},\\ \text{Prec}^{*}_{c}=\frac{\text{TP}_{c}}{\text{PP}_{c}}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}},\end{split}$ (12)
where $\text{TP}_{c}$, $\text{FP}_{c}$, and $\text{FN}_{c}$ are the true positives, false positives, and false negatives for class $c$, respectively, while $\text{P}_{c}=\text{TP}_{c}+\text{FN}_{c}$ and $\text{PP}_{c}=\text{TP}_{c}+\text{FP}_{c}$ are the real and predicted positives. In SPML, flipped positives are $\text{F}_{c}=\beta_{c}\cdot\text{P}_{c}$, where $\beta_{c}\in(0,1)$ is the proportion of flipped positives. Let $\text{PF}_{c}=\alpha_{c}\cdot\text{F}_{c}$ be the proportion of correctly classified flipped positives, where $\alpha_{c}\in[0,1]$. Thus, noisy validation metrics $\text{Rec}_{c}$ and $\text{Prec}_{c}$ are functions of their clean counterparts:
$\begin{split}\text{Rec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{P}_{c}-\text{F}_{c}}=\frac{\text{Rec}^{*}_{c}-\alpha_{c}\cdot\beta_{c}}{1-\beta_{c}},\\ \text{Prec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{PP}_{c}}=\text{Prec}^{*}_{c}\cdot\left(1-\frac{\alpha_{c}\cdot\beta_{c}}{\text{Rec}^{*}_{c}}\right).\end{split}$ (13)
These metrics depend on the label noise flip rate $\beta_{c}$ and the classifier’s behavior on flipped labels $\alpha_{c}$, which can vary, altering the noisy validation dynamic from the clean one. However, assumptions about $\alpha_{c}$ based on noise type allow us to draw some conclusions. We consider two SPML noise types: Random and Dominant. In both cases, we will demonstrate below that the expected noisy $\text{AP}_{c}$ for class $c$ relates to its true counterpart $\text{AP}_{c}^{*}$ via a class-specific scaling factor $\omega_{c}$:
$\mathbb{E}[\text{AP}_{c}]=\omega_{c}\cdot\text{AP}_{c}^{*}.$ (14)
Consequently, the expected noisy mAP, can be expressed relative to the true mAP* as follows:
$\mathbb{E}[\text{mAP}]=\frac{1}{C}\sum_{c=1}^{C}{\omega_{c}\cdot\text{AP}_{c}^{*}}=\bar{\omega}\cdot\text{mAP}^{*}+\text{Cov}\left(\omega,\text{AP}^{*}\right),$ (15)
where $\bar{\omega}=\frac{1}{C}\sum_{c=1}^{C}{\omega_{c}}$ is the mean scaling factor across classes, and $\text{Cov}\left(\omega,\text{AP}^{*}\right)$ denotes the covariance between the class-specific scaling factors and the true APs. The sign and magnitude of this covariance term determine the deviation from a simple mean-scaling relationship.
III-C1 Random SPML
In random SPML, single positive labels are chosen uniformly at random from the set of true positives for each instance. This type of noise is generally considered instance-independent [42]. While the trained model may inevitably fit this noise in the training labels, it cannot meaningfully generalize to it or predict the noisy labels in the validation set. Under this assumption, the proportion of correctly classified flipped positives, $\alpha_{c}$, on average equals the recall for that class, i.e., $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$, from which we can derive the following noisy validation metrics:
$\mathbb{E}\left[\text{Rec}_{c}\right]\approx\text{Rec}^{*}_{c},\quad\mathbb{E}\left[\text{Prec}_{c}\right]\approx\left(1-\beta_{c}\right)\cdot\text{Prec}^{*}.$ (16)
Substituting these into the AP definition, we find that the expected noisy AP for each class is scaled down by the non-flip rate:
$\mathbb{E}[\text{AP}_{c}]\approx(1-\beta_{c})\cdot\text{AP}_{c}^{*}.$ (17)
Here, the scaling factor is $\omega_{c}=(1-\beta_{c})$. Substituting this into the general mAP relationship in (15), we get the following:
$\begin{split}\mathbb{E}[\text{mAP}]&\approx\frac{1}{C}\sum_{c=1}^{C}{(1-\beta_{c}})\cdot\text{AP}_{c}^{*}<\text{mAP}^{*},\\ \mathbb{E}[\text{mAP}]&\approx(1-\bar{\beta})\cdot\text{mAP}^{*}-\text{Cov}\left(\beta,\text{AP}^{*}\right),\end{split}$ (18)
where $\bar{\beta}$ is the mean flip rate across classes. This implies that the expected noisy mAP will be lower than the true mAP*. The covariance term $\text{Cov}\left(\beta,\text{AP}^{*}\right)$ captures how the class-specific flip rates correlate with their true performances. In general, we can expect that better-performing classes experience less flipping, implying negative covariance and therefore a mitigation effect on the mean-scaling relationship in (18), i.e., $(1-\bar{\beta})\cdot\text{mAP}^{*}\leq\mathbb{E}[\text{mAP}]<\text{mAP}^{*}$.
III-C2 Dominant SPML
In contrast to random SPML, dominant SPML belongs to the category of instance-dependent label noise [42]. Here, the single-positive label is consistently the “dominant” one among the true positives. This allows the model to potentially learn this label-noise model, which can be reflected in the noisy validation mAP later in training. This results in three consecutive trends during learning [22]: i) an early learning phase where $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$ still holds, yielding similar results to random SPML; ii) a transition phase; and iii) a label-noise model fitting phase. Once the model learns the label-noise behavior, the expected proportion of correctly classified flipped positives approaches zero, as the model learns to ignore non-dominant true positives, i.e., $\mathbb{E}\left[\alpha_{c}\right]\to 0$. This implies the following expected noisy validation metrics:
$\mathbb{E}[\text{Rec}_{c}]\approx\frac{\text{Rec}_{c}^{*}}{1-\beta_{c}},\quad\mathbb{E}[\text{Prec}_{c}]\approx\text{Prec}_{c}^{*},$ (19)
which leads to the following expected AP relationship:
$\mathbb{E}[\text{AP}_{c}]\approx\frac{1}{1-\beta_{c}}\cdot\text{AP}_{c}^{*}.$ (20)
In this case, the scaling factor is $\omega_{c}={1}/(1-\beta_{c})$. The corresponding mAP relationship is as follows:
$\small\begin{split}\text{mAP}^{*}&<\mathbb{E}[\text{mAP}]=\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}\cdot\text{AP}_{c}^{*}}<1,\\ \mathbb{E}[\text{mAP}]&\approx\left(\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}}\right)\text{mAP}^{*}+\text{Cov}\left(\frac{1}{1-\beta},\text{AP}^{*}\right).\end{split}$ (21)
This indicates that the expected noisy mAP will be higher than the true $\text{mAP}_{c}^{*}$. The covariance term captures if classes with higher true APs are affected by the scaling. In general, we expect highly dominant classes to perform better than others. This leads to a negative covariance value, i.e., $\text{mAP}^{*}<\mathbb{E}[\text{mAP}]\leq\left(\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}}\right)\text{mAP}^{*}$. Consequently, noisy $\text{AP}_{c}$ are initially lower than $\text{AP}_{c}^{*}$ but later they increase their value. The noisy $\text{AP}_{c}$ approximately remain proportional to $\text{AP}_{c}^{*}$ during these phases, with the “transition phase” marking a departure from this simple scaling, possibly shifting the plateau’s appearance between them.
III-C3 Robust Plateau Detection using Teacher Model
(a) Random SPML
(b) Dominant SPML
Figure 3: True and noisy validation mAP trends of the teacher model during the training with AN on the refined BigEarthNet [14] dataset: (a) random SPML, and (b) dominant SPML.
As demonstrated in (18) and (21), although a proportional relationship generally holds between the noisy mAP and the true $\text{mAP}^{*}$, the bias term (i.e., the covariance) can introduce deviations from this relationship. However, we argue that this bias term rarely compromises the analysis, especially in the early learning stage. At the beginning, the model primarily fits correctly labelled samples [21], leading to a weak correlation between $\beta_{c}$ and $\text{AP}_{c}^{*}$. As a result, the scaling relationship predominantly governs (18) and (21). Later, when the $\text{AP}_{c}^{*}$ scores plateau, the covariance term’s contribution to the slope of the noisy mAP becomes negligible, as it exhibits little to no variation. Together, these factors enable the noisy mAP to closely mirror the true performance plateau, making it a reliable indicator for identifying the end of the early learning phase.
We propose monitoring the teacher model’s epoch-wise noisy validation mAP score to ensure a stable and reliable detection. The teacher model is maintained as an EMA of the student model’s weights. It acts as a low-pass filter, smoothing the student’s unstable learning dynamics with fluctuations caused by noisy gradients or erratic updates. As shown in Fig. 3, the noisy and true mAP scores computed from the teacher model closely follow similar plateau trends, consistent with the relationships established in (18) and (21). This smoothed behavior allows for a more accurate and robust detection of the optimal epoch for pseudo-labeling.
In practice, we online select the epoch that achieves the highest noisy validation mAP score from the teacher model, provided there is no significant improvement within the following $b_{e}$ epochs. Once this optimal epoch is detected, the model weights are automatically restored from that epoch, and pseudo-labeling is initiated concurrently with GC.
III-D Dual-EMA Pseudo-Label Generation
As discussed in Section III-B, GC relies on high-quality pseudo-labels. Since it is triggered after early learning detection on the teacher model, the teacher’s predictions are a natural choice. However, due to slow weight updates, the teacher model may exhibit a delay in incorporating recently acquired knowledge after GC begins. Therefore, we propose to combine predictions from both teacher and student models. Specifically, we apply EMA to the student’s predictions, using a smaller smoothing factor than that for the teacher model, to avoid overfitting to its own outputs while maintaining adaptability.
Let $\mathbf{p}_{i}^{T}$ and $\mathbf{p}_{i}^{S}$ denote the prediction probabilities of the teacher and student models for sample $i$, respectively. For the teacher model, we have the following:
$\mathbf{p}_{i}^{T}(e)=f(\mathbf{x}_{i};\theta^{T}(e)),$ (22)
$\theta^{T}(e)=\beta_{t}\cdot\theta^{T}(e-1)+(1-\beta_{t})\cdot\theta^{S}(e)$ (23)
with $\theta^{T}(e)$ and $\theta^{S}(e)$ being the model weights of the teacher and student models at epoch $e$, respectively. We set the smoothing coefficient $\beta_{t}=0.999$ following [43]. Then, we compute the temporarily EMA-smoothed student predictions at the $e$th iteration with the coefficient $\beta_{s}\in[0,1)$ as follows:
$\tilde{\mathbf{p}}_{i}^{S}(e)=\beta_{s}\cdot\tilde{\mathbf{p}}_{i}^{S}(e-1)+(1-\beta_{s})\cdot\mathbf{p}_{i}^{S}(e).$ (24)
We empirically set $\beta_{s}=0.8$ to guarantee the update speed of the student model’s prediction. Then, the pseudo label $\mathbf{t}_{i}(e)$ used for gradient calibration is computed as the combination of the teacher and student predictions as follows:
$\mathbf{t}_{i}=\gamma\cdot\mathbf{p}_{i}^{T}+(1-\gamma)\cdot\tilde{\mathbf{p}}_{i}^{S},$ (25)
where $e$ is omitted for simplicity. In our experiments, we set $\gamma=0.5$ to balance the contributions of the teacher and student models when generating pseudo-labels.
III-E Mixup
To further enhance robustness against label noise and improve generalization, we incorporate Mixup [15] into the GC stage. For each batch, we randomly sample a pairing input $(\mathbf{x}_{j},\mathbf{y}_{j},\mathbf{t}_{j})$ for every original input $(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{t}_{i})$, and compute:
$\tilde{\mathbf{z}}_{i}=\phi\cdot\mathbf{z}_{i}+(1-\phi)\cdot\mathbf{z}_{j}\quad\text{for}\quad\mathbf{z}\in\{\mathbf{x},\mathbf{y},\mathbf{t}\},$ (26)
where $\phi\sim\text{Beta}(\alpha,\alpha)$ and $\alpha>0$ controls the interpolation strength. The mixed samples $(\tilde{\mathbf{x}}_{i},\tilde{\mathbf{y}}_{i},\tilde{\mathbf{t}}_{i})$ are then used for training during the GC phase. This data augmentation encourages linear behavior between examples, mitigates overfitting, and stabilizes training under noisy labels.
IV Experiments
This section details the experimental setup employed to evaluate the proposed AdaGC approach against a comprehensive set of state-of-the-art (SOTA) SPML methods. We first describe the datasets utilized and the related specific SPML noise types considered. Subsequently, we outline the experimental setup, including hyperparameter setting, evaluation metrics, model architectures, and comparison methods considered in this study. Finally, we present the comparison results and an in-depth ablation study of AdaGC.
IV-A Datasets
We evaluate the proposed method on two multi-label RS benchmark datasets characterized by different modalities and spatial resolutions: refined BigEarthNet (reBEN) [44, 45, 14], derived from SAR and multispectral Sentinel-1/2 imagery with a spatial resolution of 10m, and AID-multilabel [13], collected from Google Earth aerial RGB imagery with resolutions ranging from 0.5m to 8m per pixel. reBEN focuses on land cover classification, while AID-multilabel includes both land cover categories and object-level semantic information. Due to the availability of CORINE Land Cover (CLC) [46] pixel-wise masks for reBEN, we were able to simulate the two types of label noise described in Section III-C. In contrast, AID’s high spatial resolution introduces greater semantic complexity, making it challenging to determine a dominant class. Therefore, we simulate only Random single-positive label noise on AID.
IV-A1 Refined BigEarthNet
It contains 549,488 Sentinel-1/2 image pairs of size of $120\times 120$ pixels covering 10 European countries, each aligned with a pixel-level CLC reference map. Multi-label annotations are derived from these maps, resulting in 19 land cover classes (see Table I). Following the splits in [14], we use a 2:1:1 ratio for training, validation, and testing. Each input has 14 channels by concatenating 2 Sentinel-1 and 12 Sentinel-2 L2A bands. Leveraging the reference maps, on this dataset we simulate both types of SPML label noise (Random and Dominant). Dataset statistics for the two cases are provided in Table I.
TABLE I: Refined BigEarthNet Land Cover Classes and Related Flip
Rates $\beta_{c}$ on the Training Subset in the Two Cases of
Random and Dominant SPML Label Noise. Macro
and Micro Averages Are Provided, as Well as
the Average Number of True Positive
Labels per Image
Class name Random $\beta_{c}$ Dominant $\beta_{c}$ Support
Urban fabric 0.73 0.84 33626
Industrial or commercial units 0.72 0.91 6398
Arable land 0.63 0.51 88236
Permanent crops 0.70 0.74 15190
Pastures 0.66 0.62 42968
Complex cultivation patterns 0.73 0.77 51034
Land principally occupied by agriculture, with significant areas of natural vegetation 0.73 0.86 60127
Agro-forestry areas 0.64 0.54 15082
Broad-leaved forest 0.69 0.65 64164
Coniferous forest 0.66 0.54 76344
Mixed forest 0.70 0.67 78791
Natural grassland and sparsely vegetated areas 0.69 0.82 6984
Moors, heathland and sclerophyllous vegetation 0.70 0.70 6329
Transitional woodland, shrub 0.70 0.85 64198
Beaches, dunes, sands 0.72 0.94 738
Inland wetlands 0.67 0.78 11344
Coastal wetlands 0.70 0.70 670
Inland waters 0.67 0.73 30466
Marine waters 0.08 0.05 35914
Micro average 0.65 0.65
Macro average $\bar{\beta}$ 0.66 0.70
Average #labels/image 2.9
IV-A2 AID-multilabel
It extends the AID scene classification dataset [47] with multi-label annotations. The dataset comprises 3,000 high-resolution RGB aerial image patches ($600\times 600$ pixels) across 17 classes, collected from China, U.S., and several European countries. We exclude the extremely rare “mobile home” class (only 2 samples), resulting in 16 classes. Following [13], 20% of the data is held out for testing. We then use 25% of the remaining training set for validation. Dataset statistics are provided in Table II. For simplicity, in the following we use AID to refer to AID-multilabel.
TABLE II: AID Multilabel Classes and Related Flip Rates $\beta_{c}$ on the
Training Subset in the Case of Random SPML Label Noise.
Macro and Micro Averages Are Also Provided, as Well
as the Average Number of True Positive
Labels per Image
Class name Random $\beta_{c}$ Support
Airplane 0.83 59
Bare-soil 0.80 883
Buildings 0.82 1313
Cars 0.84 1211
Chaparral 0.72 58
Court 0.85 207
Dock 0.91 163
Field 0.74 129
Grass 0.79 1366
Pavement 0.81 1408
Sand 0.47 153
Sea 0.77 133
Ship 0.86 176
Tanks 0.83 65
Trees 0.78 1441
Water 0.84 509
Micro average 0.81
Macro average $\bar{\beta}$ 0.79
Average #labels/image 5.2
IV-B Experimental Setup
We adopt ResNet-34 and ResNet-50 as backbone networks for the reBEN and AID datasets, respectively, with batch sizes of 512 and 32. We choose ResNet-34 for reBEN as it has fewer parameters and lower computational complexity, which significantly reduces the overall runtime of our experiments. Despite being lighter, it shares a similar architecture with ResNet-50 and generally achieves comparable performance. Moreover, since our primary goal is to evaluate the effectiveness of the proposed method relative to baseline settings, rather than to maximize absolute accuracy, ResNet-34 provides a suitable and efficient compromise. Given the AID dataset’s limited size and RGB data, we initialize the network with ImageNet-pretrained weights. In this setup, we consider the realistic scenario of absence of clean labels for training and hyperparameter tuning. Hence, all methods’ hyperparameters were tuned via grid search by evaluating their mAP score on the noisy validation set. To manage the computational burden associated with the large number of methods and hyperparameters, we performed this tuning process using a subset of the respective datasets (20% of training/validation sets for reBEN and 50% for AID) and for a reduced number of epochs (10 epochs instead of 70 as in the final runs). For AdaGC, we set the Mixup Beta distribution parameter $\alpha$ to 1, and the GC loss weight $\lambda$ to 3, 5, and 8 for reBEN-Random, reBEN-Dominant, and AID, respectively, selected via hyperparameters tuning.
For the final evaluation, we utilized the optimal hyperparameters identified during the tuning phase. Each model was then trained on the full training dataset for a total of 70 epochs, after which each model was evaluated on the test dataset. For the proposed AdaGC approach, the noisy validation mAP from the teacher model, as discussed in Section III-C, was used to detect the optimal epoch for initiating GC with pseudo-label generation. To ensure the robustness and statistical reliability of our results, each method underwent three Monte Carlo runs, allowing us to quantify the variability in performance due to random weight initialization and dataset shuffling. We report the average values of all metrics along with their standard deviations (computed using an unbiased estimator).
For comprehensive evaluation, in addition to mAP, we report two ranking-based metrics, i.e., coverage and ranking loss (Rankloss), and four accuracy-based metrics, i.e., overall accuracy (OA), mean F1 score (mF1), mean precision (mprecision), and mean recall (mrecall). The metrics are defined as follows.
  • Coverage measures how far we need to go down the ranked label list to cover all true labels:
    $\frac{1}{N}\sum_{i=1}^{N}\max_{y\in\mathbf{y}_{i}}\text{rank}_{f}(\mathbf{x}_{i},y)-1,$ (27)
    where $\text{rank}_{f}(x_{i},y)$ is the position of the ground-truth (GT) label $y_{i,c}$ in the ranking list predicted by $f$;
  • Ranking loss computes the average fraction of label pairs that are incorrectly ordered:
    $\small\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathbf{y}_{i}||\overline{\mathbf{y}_{i}}|}\sum_{(y_{1},y_{0})\in\mathbf{y}_{i}\times\overline{\mathbf{y}_{i}}}\mathbb{I}(f(\mathbf{x}_{i},y_{1})\leq f(\mathbf{x}_{i},y_{0})),$ (28)
    where $\mathbf{y}_{i}$ and $\overline{\mathbf{y}_{i}}$ are the sets of relevant (positive) and irrelevant (negative) labels, respectively, and $\mathbb{I}(\texttt{True})=1$ otherwise 0;
  • OA measures the percentage of samples with all labels correctly predicted:
    $\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}\mathbb{I}(\hat{\mathbf{y}}_{i,c}=\mathbf{y}_{i,c});$ (29)
  • mF1 is the average of F1-scores computed per class:
    $\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Prec}_{c}\cdot\text{Rec}_{c}}{\text{Prec}_{c}+\text{Rec}_{c}}.$ (30)
We compared AdaGC against a wide array of existing methods, categorized as follows.
  • Baseline methods: We considered several baselines [12], such as AN, AN-LS, Weak (WAN), and EPR. All use a BCE loss at their core, while EPR uses only the BCE positive term and adds a regularization term, assuming the a-priori knowledge of the average number of positive labels per sample (see Tables I and II).
  • General label noise methods: From this category, we considered the ELR approach adapted as in [10], the LL family of approaches (i.e., LL-R, LL-Ct, and $L$-Cp) [16], and the BoostLU [19] in combination with $L$-Ct.
  • PUL methods: We considered a single SOTA method from this category, i.e., Dist-PU [39], adapted for SPML by simply applying it to each class independently.
  • SPML-specific methods: The main comparison methods we considered for SPML are ROLE [12], GR Loss [18], LAGC [17], and Mutual label enhancement for sIngle-positive Multi-label lEarning (MIME) [20].
  • Upper bound methods: For analytical comparison with GT training, we considered two cases: i) training with GT multi-labels using a BCE loss (GT w/ BCE), and ii) the with true Negatives (IUN) approach [12], where in addition to the single positive labels, all true negatives are provided for training. IUN allows us to assess the performance in a hypothetical scenario where all true negatives have been retrieved. False negatives are ignored during training, meaning that the model is still provided with a single positive label for each image.
IV-C Results
TABLE III: Test Performance Comparison of Different Methods on the reBEN-Random Dataset. The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic. The Related Standard Deviations Are Reported in Brackets
mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
Ground-truth Training GT w/ BCE 71.75 (0.04) 4.264 (0.003) 3.428 (0.005) 93.09 (0.01) 66.63 (0.12) 72.07 (0.02) 63.24 (0.35)
IUN 70.46 (0.11) 4.331 (0.009) 3.608 (0.024) 91.98 (0.01) 56.00 (0.19) 84.75 (0.55) 44.66 (0.18)
Baseline Methods AN 54.97 (0.12) 4.849 (0.002) 5.847 (0.003) 87.13 (0.04) 28.15 (0.30) 72.93 (1.67) 19.00 (0.21)
AN-LS 54.40 (0.66) 5.375 (0.114) 7.051 (0.309) 86.90 (0.03) 25.47 (0.44) 75.90 (0.73) 17.07 (0.31)
WAN 59.43 (0.34) 4.660 (0.010) 5.089 (0.037) 90.16 (0.05) 60.15 (0.20) 55.96 (0.36) 66.14 (0.65)
EPRa 50.74 (0.40) 5.206 (0.022) 6.825 (0.053) 80.13 (0.08) 51.22 (0.11) 38.25 (0.13) 85.59 (0.19)
General Label Noise Methods ELR 62.40 (0.08) 4.860 (0.021) 4.890 (0.021) 91.45 (0.04) 63.97 (0.07) 62.63 (0.36) 68.26 (0.25)
$L$-Cp 57.27 (0.13) 4.983 (0.013) 6.068 (0.038) 89.60 (0.03) 49.35 (0.58) 65.47 (0.40) 43.67 (0.88)
$L$-Ct 62.35 (0.23) 4.722 (0.010) 4.856 (0.029) 91.76 (0.03) 54.87 (0.35) 76.27 (1.09) 46.41 (0.18)
$L$-R 59.70 (0.27) 4.688 (0.012) 5.054 (0.042) 90.85 (0.05) 53.91 (0.56) 70.44 (1.08) 45.45 (0.45)
BoostLU 61.77 (0.63) 4.662 (0.011) 4.780 (0.023) 91.59 (0.01) 55.41 (0.61) 74.80 (1.30) 46.32 (0.36)
PUL Methods Dist-PUa 61.98 (0.26) 4.860 (0.016) 5.111 (0.024) 90.67 (0.04) 61.49 (0.35) 59.11 (0.47) 68.83 (0.52)
SPML Methods ROLE 55.73 (0.36) 6.479 (0.007) 6.424 (0.020) 91.18 (0.05) 59.65 (0.31) 61.90 (0.45) 58.36 (0.35)
GR 57.38 (0.49) 4.914 (0.034) 5.855 (0.120) 89.77 (0.08) 56.79 (0.85) 57.91 (0.16) 56.06 (1.56)
LAC 57.88 (0.30) 4.801 (0.021) 5.384 (0.036) 87.33 (0.25) 28.97 (1.38) 72.29 (0.62) 19.85 (1.11)
MIME 58.64 (0.35) 4.894 (0.017) 5.632 (0.064) 90.66 (0.13) 55.13 (0.36) 62.40 (0.46) 50.39 (0.21)
Proposed AdaGC 68.18 (0.35) 4.513 (0.012) 4.073 (0.033) 92.52 (0.02) 64.65 (0.37) 70.36 (0.84) 63.37 (0.82)
a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
TABLE IV: Test Performance Comparison of Different Methods on the reBEN-Dominant Dataset. The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic. The Related Standard Deviations Are Reported in Brackets
mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
Ground-truth Training GT w/ BCE 71.75 (0.04) 4.264 (0.003) 3.428 (0.005) 93.09 (0.01) 66.63 (0.12) 72.07 (0.02) 63.24 (0.35)
IUN 64.46 (0.07) 5.289 (0.013) 5.908 (0.007) 90.78 (0.01) 47.38 (0.09) 86.11 (0.95) 35.73 (0.06)
Baseline Methods AN 55.91 (0.29) 6.533 (0.075) 8.128 (0.084) 88.70 (0.01) 34.25 (0.26) 84.35 (0.49) 24.08 (0.15)
AN-LS 53.65 (0.12) 8.474 (0.088) 14.878 (0.282) 88.61 (0.00) 33.93 (0.16) 85.99 (0.99) 23.75 (0.12)
WAN 58.69 (0.28) 6.627 (0.054) 8.153 (0.044) 90.76 (0.02) 50.56 (0.27) 74.22 (1.12) 41.50 (0.14)
EPRa 49.15 (0.36) 6.150 (0.034) 9.003 (0.107) 87.22 (0.07) 53.34 (0.20) 48.48 (0.47) 64.38 (0.39)
General Label Noise Methods ELR 54.48 (0.33) 9.149 (0.015) 11.429 (0.087) 91.02 (0.02) 50.08 (0.10) 76.38 (0.81) 42.05 (0.13)
$L$-Cp 47.70 (0.20) 6.061 (0.016) 8.968 (0.012) 88.04 (0.04) 53.17 (0.11) 51.47 (0.49) 59.48 (0.30)
$L$-Ct 54.11 (0.27) 6.287 (0.107) 8.014 (0.216) 90.88 (0.14) 53.67 (0.12) 70.67 (0.90) 47.82 (0.25)
$L$-R 54.41 (0.13) 7.369 (0.081) 8.662 (0.055) 91.01 (0.02) 53.59 (0.20) 70.83 (0.96) 46.88 (0.16)
BoostLU 47.36 (0.34) 6.711 (0.110) 9.767 (0.376) 88.62 (0.21) 54.70 (0.45) 57.40 (1.38) 58.96 (0.46)
PUL Methods Dist-PUa 53.51 (0.32) 6.224 (0.109) 8.770 (0.260) 88.23 (0.09) 53.17 (0.29) 56.21 (1.78) 60.53 (0.17)
SPML Methods ROLE 48.54 (0.41) 12.465 (0.083) 14.591 (0.157) 90.81 (0.03) 50.24 (0.46) 75.38 (1.05) 40.88 (0.47)
GR 55.73 (0.23) 5.765 (0.050) 7.238 (0.097) 89.20 (0.00) 38.28 (0.18) 81.43 (0.56) 27.57 (0.16)
LAC 54.67 (0.56) 6.029 (0.074) 8.170 (0.206) 88.57 (0.02) 33.43 (0.63) 80.57 (1.58) 23.67 (0.37)
MIME 55.61 (0.07) 6.417 (0.061) 8.982 (0.125) 89.90 (0.03) 43.54 (0.39) 76.38 (1.19) 33.36 (0.55)
Proposed AdaGC 60.22 (0.24) 5.618 (0.004) 6.926 (0.016) 90.15 (0.05) 58.78 (0.44) 61.28 (0.98) 62.48 (0.62)
a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
This section presents a detailed analysis of the experimental results reported in Tables III, IV, and V, comparing the performance of AdaGC against various baselines, PUL, SPML-specific, and general label noise methods across the reBEN and AID datasets under different SPML noise types. Several general trends are observable across all experimental scenarios. Our proposed AdaGC consistently achieves the best mAP values across all three evaluated scenarios. Furthermore, it demonstrates SOTA SPML performance in terms of coverage, rankloss, and mF1, highlighting its overall effectiveness and robustness in handling SPML learning for RS imagery. The upper-bound method IUN, which leverages knowledge of true negatives while only observing single positives, consistently achieves a good mAP. However, due to its bias towards single positive labels, it typically exhibits high mprecision but comparatively lower mrecall, which consequently impacts its mF1. The AN and AN-LS baselines are particularly sensitive to the inherent bias introduced by single positive labels and the prevalence of false negatives. This sensitivity is reflected in their high mprecision but significantly low mrecall, leading to a generally low mAP across all datasets. In contrast, the WAN method demonstrates greater robustness compared to AN and AN-LS. It achieves a better balance between mprecision and mrecall, making it the best-performing baseline method in terms of mAP across the considered scenarios.
EPR employs a distinct strategy by training solely on single positive samples and incorporating a regularization term to enforce the prediction of an average number of positive labels. This approach consistently results in EPR being the model with the highest mrecall across all datasets and noise types. However, given its reliance only on positive supervision, its mprecision is generally lower compared to methods that attempt to identify true negatives. Additionally, the consistently high mrecall scores of EPR are attributable to its unique reliance on a-priori knowledge of the average number of positive labels per image, a piece of information not utilized by other methods under comparison. Next subsections discuss the results obtained on each specific dataset and scenario.
IV-C1 reBEN-Random
TABLE V: Test Performance Comparison of Different Methods on the AID Dataset. The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic. The Related Standard Deviations Are Reported in Brackets
mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
Ground-truth Training GT w/ BCE 88.90 (0.09) 5.784 (0.036) 1.649 (0.177) 94.35 (0.14) 83.70 (0.34) 87.43 (1.02) 81.91 (0.63)
IUN 82.85 (0.33) 6.311 (0.060) 2.741 (0.087) 91.21 (0.37) 71.17 (2.28) 95.04 (1.41) 60.64 (2.15)
Baseline Methods AN 65.41 (1.28) 8.403 (0.177) 10.908 (0.349) 71.74 (0.19) 27.16 (0.12) 86.42 (2.26) 16.92 (0.09)
AN-LS 64.89 (0.71) 11.175 (0.189) 21.113 (1.477) 70.81 (0.25) 20.11 (0.57) 80.53 (5.82) 12.22 (0.44)
WAN 68.70 (0.49) 7.481 (0.185) 7.575 (0.647) 78.36 (0.17) 50.14 (0.47) 84.57 (1.09) 36.95 (0.35)
EPRa 68.62 (1.87) 7.277 (0.093) 8.453 (0.390) 84.91 (0.57) 70.09 (1.12) 62.33 (0.85) 81.76 (1.78)
General Label Noise Methods ELR 69.97 (0.69) 7.176 (0.014) 5.401 (0.052) 88.25 (0.20) 59.36 (0.89) 74.95 (0.96) 56.19 (0.35)
$L$-Cp 64.71 (1.07) 7.629 (0.132) 7.426 (0.465) 78.73 (0.57) 45.64 (1.16) 76.96 (3.89) 36.35 (0.72)
$L$-Ct 67.48 (1.53) 8.302 (0.260) 6.846 (0.374) 85.78 (0.65) 55.78 (0.91) 78.65 (4.73) 48.19 (0.88)
$L$-R 69.56 (0.43) 8.314 (0.023) 7.958 (0.313) 80.75 (0.58) 51.42 (2.25) 85.28 (2.87) 39.15 (2.38)
BoostLU 68.99 (1.11) 7.218 (0.113) 5.733 (0.308) 85.79 (0.94) 54.86 (0.83) 79.39 (1.79) 47.40 (0.97)
PUL Methods Dist-PUa 71.38 (0.28) 6.493 (0.112) 3.761 (0.177) 90.02 (0.11) 60.59 (3.33) 78.60 (6.94) 57.09 (3.12)
SPML Methods ROLE 56.68 (0.79) 13.238 (0.053) 19.038 (0.294) 76.44 (0.38) 43.66 (0.64) 84.18 (3.33) 31.10 (0.49)
GR 62.69 (0.79) 11.053 (0.091) 20.911 (0.577) 72.34 (0.08) 28.62 (0.57) 83.20 (4.34) 18.36 (0.31)
LAC 62.61 (1.01) 7.746 (0.086) 9.187 (0.231) 72.74 (0.03) 28.84 (1.87) 82.41 (4.61) 18.56 (1.31)
MIME 61.96 (1.94) 7.150 (0.166) 7.016 (0.799) 73.94 (0.28) 31.96 (1.03) 81.02 (2.73) 20.82 (0.66)
Proposed AdaGC 74.82 (0.90) 6.607 (0.054) 4.381 (0.154) 86.51 (0.49) 60.68 (2.24) 77.86 (1.33) 60.03 (2.15)
a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
Table III summarizes the performance of all methods on the reBEN dataset when subjected to Random SPML noise. Our proposed AdaGC demonstrates superior overall performance, achieving the best mAP, OA, mF1, coverage, and rankloss scores. While it does not achieve the best performance in mprecision and mrecall, AdaGC shows the best balance between two, as reflected in the mF1 score.
General label noise and PUL methods show good performance in this scenario, outperforming SPML-specific methods. Additionally, WAN demonstrates to be a strong baseline, outperforming all other baselines as well as many other comparison methods. The reason might be found in the hyperparameter selection. Indeed, we tuned them for each method individually by monitoring the noisy validation mAP. While noisy mAP can be a good indicator for early learning detection when evaluated continuously during training, its fluctuations in hyperparameters selection, combined with methodologies that stem from AN-like approaches (e.g., pseudo-labelling or label correction approaches), can result in suboptimal hyperparameters’ selection. However, we argue that methods benchmarking should reflect real-world settings as much as possible, which is ensured in our experimental setting.
Comparing the results with GT training, we can observe that the AdaGC approaches the performance of training with GT multi-labels, especially in terms of mF1. Compared to IUN, AdaGC achieves a worse mAP score and a better mF1 score. Although false negatives do not affect IUN in mAP results, the mF1 scores emphasize the importance of methodologies that leverage pseudo-labels or label correction. Such approaches are vital for accurately reconstructing multi-labels, thus preventing classifier confidence bias and achieving higher mF1 scores.
IV-C2 reBEN-Dominant
While Random SPML can be a good simulation for cases where both real-world single labels and GT multi-labels are difficult to gather, it provides a simplified scenario. Indeed, random single labels can still provide all the necessary information for describing the land cover, since the single labels are chosen randomly among the GT labels. On the contrary, dominant single labels are completely driven by the actual observed image, leading to cases where specific non-dominant class patterns are entirely missing from the training set, making this scenario more challenging. This can be observed in Table IV, where all methods show lower absolute performance compared to their counterparts in Table III.
Despite this increased difficulty, the proposed AdaGC maintains its leading position, achieving the highest mAP, mF1, coverage, and rankloss scores. Instead, the comparison methods struggle to outperform the WAN and EPR baseline methods. WAN shows the second best mAP score, and its mF1 score surpasses all SPML-specific methods. Only methods based on the LL family and Dist-PU are able to outperform WAN in terms of mF1, struggling against EPR. Due to the regularization term that affects the number of predicted positive labels, EPR can overcome the strong bias introduced by Dominant SPML in terms of mF1 score, despite its low mAP score.
IV-C3 AID
(a) reBEN
(b) AID
Figure 4: Histograms of the number of GT labels per image in the training sets.
Table V displays the performance of all methods on the AID dataset with Random SPML noise. This dataset, benefiting from ImageNet pre-training and richer spatial information, generally shows higher absolute performance metrics across the board compared to reBEN, despite its generally higher flip rates $\beta_{c}$. Consistent with its overall very good performance, AdaGC emerges at the top of the ranking for most metrics. It achieves the best mAP score, and the second best mF1, mrecall, coverage, and rankloss scores. Notably, Dist-PU shows outstanding performance, on par with AdaGC’s mF1, and achieves the best OA, coverage, and rankloss scores. Additionally, EPR shows the best mF1 score. Nevertheless, the top performance of Dist-PU and EPR can be explained in terms of prior information available. Indeed, both methods exploit some a-priori information regarding the GT multi-labels, which is not the case for the proposed AdaGC. Dist-PU assumes that the knowledge of the prior distributions of all classes is available, and exploits this information to match it with the class priors of the predicted labels, while EPR uses the knowledge of the expected number of positive labels per image. This information is particularly effective on the AID dataset, as it shows, in general, a higher number of labels for each image, as shown in Fig. 4. This makes EPR exhibit the best mrecall among all methods, on par with GT training, which is high enough to counterbalance the worst mprecision across the board, which highlights a label overestimation tendency.
IV-D Ablation Experiments
We conduct ablation experiments on reBEN under both types of label noise to evaluate the effectiveness of individual components in AdaGC. Without making specific claims, we report the accuracies of the teacher models below.
IV-D1 Early Learning Detection
(a) Random-mAP
(b) Random-mF1
(c) Dominant-mAP
(d) Dominant-mF1
Figure 5: Validation accuracies (mAP and mF1) versus training time (epoch) obtained with predefined warm-up length (10, 20, 30) and our proposed early learning detection strategy (AdaGC).
We evaluate the effectiveness of early learning detection for adaptively triggering GC by comparing AdaGC with variants that use manually predefined trigger points. Fig. 5 plots their true validation accuracies (with respect to the GT labels) over training time. As shown, the proposed adaptive detection strategy successfully activates GC at an appropriate point, resulting in the most significant performance improvements. In contrast, triggering GC too early or too late leads to suboptimal results. Notably, activating GC too late leads to even worse performance, highlighting the detrimental effect of overfitting to label noise. Furthermore, Table VI reports the final test accuracies. AdaGC achieves competitive or comparable mAP and mF1 scores compared to those obtained by using fixed warm-up lengths, indicating the robustness of the proposed early learning detection strategy.
TABLE VI: Accuracies (%) on the Test Set Obtained With Different GC Trigger Settings After 70 Epochs
Random Dominant
mAP $\uparrow$ mF1 $\uparrow$ mAP $\uparrow$ mF1 $\uparrow$
AdaGC (after $\sim$17 epochs) 68.18 64.65 60.22 58.78
Trigger after 10 epochs 65.78 64.06 57.16 58.58
Trigger after 20 epochs 65.87 62.39 59.35 59.28
Trigger after 30 epochs 63.18 54.77 58.62 55.71
IV-D2 Teacher and Student Models
We examine the roles of the student and teacher models in two modules, i.e., early learning detection and pseudo-label generation.
We first illustrate the smoothness of the noisy validation accuracies produced by the student and teacher models in Fig. 6. The accuracy curves from both models exhibit similar trends. Nevertheless, those from the teacher models are noticeably smoother than those from the student models. This makes teacher models more suitable for early learning detection. In contrast, using student models for detection is unstable, as the detection is more likely to fall into a local maximum, leading to suboptimal results.
(a) Random
(b) Dominant
Figure 6: Noisy validation mAP (with respect to noisy labels) versus training time (epoch) obtained by AN on the reBEN dataset.
TABLE VII: Accuracies (%) on the Test Set Obtained With Pseudo Labels
Generated by Different Models, Where T and S Represent
Teacher and Student Models, Respectively
Random Dominant
mAP $\uparrow$ mF1 $\uparrow$ mAP $\uparrow$ mF1 $\uparrow$
AdaGC (T+S, $\gamma=0.5$) 68.18 64.65 60.22 58.78
Only T ($\gamma=1$) 61.74 63.49 42.01 46.40
Only S ($\gamma=0$) 67.93 61.52 61.11 57.42
T+S (S w/o EMA, $\gamma=0.5$) 53.78 61.94 27.50 35.71
We then investigate the roles of the student and teacher models in pseudo-label generation. As shown in Table VII, combining predictions from both models yields the most robust performance. In the Random noise setting, relying solely on either the student or the teacher model fails to achieve optimal mAP and mF1 scores. In the Dominant noise setting, using only the student model produces more stable results than using only the teacher model. This is partly due to the teacher model’s relatively slow update rate, which lags behind the student model’s more responsive prediction updates. As a result, the teacher may fail to adapt quickly enough to changing training dynamics, making it more vulnerable to overfitting to the dominant noisy labels. In such scenarios, swift and timely pseudo-label updates are essential for maintaining performance and robustness. Nevertheless, combining predictions from both models offers a balance between stability and adaptability, ensuring robust performance across different types of label noise.
Finally, we evaluate the role of applying EMA to student predictions by replacing them with the student model’s direct online predictions. The results, reported in the last row of Table VII, show a significant performance drop in both cases, with a more sever decrease in the Dominant setting. This highlights the importance of using EMA to stabilize the student predictions and ensure higher-quality pseudo-labels in GC.
IV-D3 Role of Mixup
We remove Mixup from AdaGC and report the results in Table VIII to assess its impact. As shown, the absence of Mixup leads to a consistent drop in performance across both noise settings, particularly in the Random case, where mAP and mF1 decrease by approximately 2%. These findings suggest that Mixup enhances generalization, especially under random noise, by encouraging smoother decision boundaries. Although its effect is less pronounced under Dominant noise, it still provides measurable benefits, underscoring its complementary role in improving the robustness of AdaGC.
TABLE VIII: Accuracies (%) on the Test Set Obtained With and Without Mixup
Random Dominant
mAP $\uparrow$ mF1 $\uparrow$ mAP $\uparrow$ mF1 $\uparrow$
AdaGC 68.18 64.65 60.22 58.78
w/o Mixup 66.00 62.08 59.05 59.19
V Conclusion
In this work, we have presented AdaGC, a novel and generalizable framework for SPML tailored to RS imagery, which addresses overfitting to false negatives via a gradient calibration mechanism. To ensure its operational effectiveness, we introduce a simple yet theoretically grounded indicator to adaptively activate GC when it detects the end of early learning during an initial warm-up stage. In addition, a dual-EMA module is introduced to produce robust pseudo-labels that support the GC process, along with Mixup to further boost the robustness of training.
Extensive experiments on two benchmark RS datasets under two single-positive realistic label noise simulation settings, Random and Dominant, demonstrate the effectiveness and robustness of AdaGC. AdaGC consistently outperforms existing SPML baselines across different scenarios. Our work fills a critical gap in the literature, where SPML remains largely unexplored in the RS domain. While AdaGC performs well in most settings, it shows relative limitations under the dominant single-positive label noise condition. As this scenario is common in many real-world annotation practices, future work will focus on enhancing AdaGC’s performance under such conditions, further advancing scalable and reliable multi-label learning in RS.

### v2

@@
 To bridge this gap, we propose Adaptive Gradient Calibration (AdaGC), a novel and generalizable SPML framework tailored to RS imagery.
-AdaGC adopts a gradient calibration (GC) mechanism combined with Mixup and a dual exponential moving average (EMA) module for robust pseudo-label generation.
-To maximize AdaGC’s effectiveness, we introduce a simple yet theoretically grounded indicator to adaptively trigger GC after an initial warm-up stage based on training dynamics, thereby guaranteeing the effectiveness of GC in mitigating overfitting to label noise.
+AdaGC adopts a gradient calibration (GC) mechanism with a dual exponential moving average (EMA) module for robust pseudo-label generation.
+We introduce a theoretically grounded, training-dynamics-based indicator to adaptively trigger GC, which ensures GC’s effectiveness by preventing it from being affected by model underfitting or overfitting to label noise.
 Extensive experiments on two benchmark RS datasets under two distinct label noise types demonstrate that AdaGC achieves state-of-the-art (SOTA) performance while maintaining strong robustness across diverse settings.
+The codes and data will be released at https://github.com/rslab-unitrento/AdaGC.
 I Introduction
-With the rapid development of satellite missions, vast amounts of remote sensing (RS) imagery have become increasingly accessible, serving as a critical data source for numerous applications [1, 2].
-As a fundamental task, image classification plays a vital role in connecting imagery with downstream applications in the RS domain [3, 4, 5].
-However, most existing RS studies adopt the single-label classification (SLC) paradigm, assigning a single semantic label to each image [6, 7], which oversimplifies the inherent complexity of real-world RS scenes [8, 9].
-In practice, RS images usually contain diverse land cover types, objects, or functional zones within a single scene.
+With the rapid development of satellite missions, vast amounts of remote sensing (RS) imagery have become increasingly accessible, serving as a critical data source for numerous applications [zhao_artificial_2024, zhu_foundations_2024].
+Image classification [mehmood_remote_2022, dimitrovski_current_2023, he_recent_2018], a fundamental task to connect RS imagery with downstream applications, still largely adopts the single-label classification (SLC) paradigm.
+However, assigning a single label to each image oversimplifies the complexity of real-world RS scenes [Hua_2019_JPRS, Sumbul_2020_ACCESS].
+In practice, RS images often contain diverse land cover types, objects, or functional zones.
 As illustrated in Fig.
-1, very-high resolution (VHR) imagery often captures land use and land cover information as well as fine-grained objects such as buildings, vehicles, and vegetation within a limited area.
-Lower-resolution imagery often spans heterogeneous regions, such as agricultural mosaics or urban–rural transitions.
-In both cases, multi-label classification (MLC) offers a more informative, comprehensive, and realistic semantic representation of RS imagery compared to traditional SLC methods.
-However, MLC faces a major challenge in acquiring complete and accurate annotations due to the difficulty and high cost in the annotation process [10].
-This challenge is particularly critical in the deep learning (DL) era, where large-scale labelled datasets are essential [11].
-A practical compromise is to annotate each image with only one relevant label, even when multiple classes are present, which leads to the single-positive multi-label learning (SPML) setting [12].
+1, both VHR and lower-resolution imagery commonly capture heterogeneous content, making multi-label classification (MLC) a more informative representation than SLC.
+A key challenge for MLC is the difficulty and cost of acquiring complete, accurate annotations [Burgert_2022_TGRS], especially in the deep learning (DL) era where large-scale labeled datasets are crucial [zhu_deep_2017].
+A practical workaround is to annotate only one relevant label per image, leading to the single-positive multi-label learning (SPML) setting [Cole_2021_CVPR].
 In this setting, the model is weakly supervised with one positive label per image, yet it is expected to recover the complete set of relevant labels.
-Supervision in SPML is inherently weak, characterized by missing labels.
-In practice, even expert annotators often miss relevant labels in multi-label settings, making the false negative issue pervasive in MLC.
-SPML can thus be viewed as an extreme case of this problem and also offers a promising solution to completing label sets in single-label annotation scenarios.
+The supervision in SPML is characterized by missing labels.
+In practice, even expert annotators often miss relevant labels, making the false negative issue pervasive in MLC.
+SPML can thus be viewed as an extreme case of this problem.
 Refer to caption Refer to caption Refer to caption
 (a) Single-label: court Multi labels: buildings, cars, pavement, court, grass, trees (b) Single-label: arable land (dominant) Multi-labels: arable land, urban fabric, inland waters, complex cultivation patterns
-Figure 1: Single- and multi-label annotation examples from (a) AID-multilabel [13] and (b) refined BigEarthNet [14] datasets, where the corresponding CLC mask is presented for reference.
+Figure 1: Single- and multi-label annotation examples from (a) AID-multilabel [hua_relation_2020] and (b) refined BigEarthNet [clasen_reben_2025_arxivonly] datasets, where the corresponding CLC mask is presented for reference.
 Compared to the single-class labels, the multi-label annotations can more comprehensively describe the scene.
-Although SPML is promising in scalability and annotation efficiency, it also introduces significant ambiguity that can bias the learning process.
-A common baseline approach is the AN strategy [12], which treats all unobserved labels as negatives, unavoidably introducing bias into learning.
-In the computer vision (CV) domain, several strategies have been proposed to address its limitations, such as noise-robust regularization terms, online label-correction/pseudo-labeling methods, and interpretability-driven frameworks [12, 15, 16, 17, 18, 19, 20].
+Although SPML is promising in scalability and annotation efficiency, it introduces significant ambiguity.
+A common baseline approach is the AN strategy [Cole_2021_CVPR], which treats all unobserved labels as negatives, unavoidably introducing bias into learning.
+In the computer vision (CV) domain, several strategies have been proposed to address its limitations, such as noise-robust regularization terms, online label-correction/pseudo-labeling methods, and interpretability-driven frameworks [Cole_2021_CVPR, Zhang_2018_ICLR, Kim_2022_CVPR, Xie_2022_NIPS, Chen_2024_IJCAI, Kim_2023_CVPR, Liu_2023_ICML].
 Despite the advancements in CV, the SPML problem remains underexplored in RS.
-SPML is particularly valuable for RS image classification, where annotation typically requires greater domain expertise and incurs higher costs compared to natural image datasets.
+SPML is particularly valuable for RS image classification, where annotation typically requires greater domain expertise and incurs higher costs.
 These challenges make the application of SPML in RS more necessary and also more difficult.
 When directly applying existing SPML methods to RS data, several challenges arise.
-One major issue is the reliability of pseudo-labeling methods.
-VHR RS images often include small, sparsely distributed targets, while medium-to-low-resolution imagery tends to capture broader regional semantics.
+One major issue is the reliability of pseudo-labels.
+very-high resolution (VHR) RS images often include small, sparsely distributed targets, while medium-to-low-resolution imagery tends to capture broader regional semantics.
 Both cases can easily result in low inter-class separability and high intra-class variability, hindering the effectiveness of simple pseudo-labeling strategies such as confidence thresholding.
 Furthermore, the distinct characteristics of MLC versus multi-class classification (MCC) suggest that commonly used learning-from-noisy-labels strategies, which are often designed for MCC, should be carefully tailored to SPML.
-Moreover, although early learning theory indicates that models gradually reach a peak generalization performance in the early stage of training before overfitting to label noise [21, 22], most existing SPML methods rely on predefined warm-up schedules and overlook this dynamic behaviour.
-This, to some extent, limits the model’s capability in pseudo-labeling.
+Moreover, most existing pseudo-label-based SPML methods depend on fixed warm-up schedules, which often result in suboptimal performance due to model underfitting or overfitting to label noise.
 In this work, we propose a novel and generalizable framework for SPML in RS, termed Adaptive Gradient Calibration (AdaGC), to address the aforementioned challenges.
 The framework consists of three key components.
-First, we introduce a dual exponential moving average (EMA) module for robust pseudo-label estimation, which leverages the consensus between the temporally smoothed predictions of the student model via EMA and the outputs of the teacher model, whose weights are updated via EMA of the student’s parameters.
-Second, the generated pseudo-labels are used to guide a Gradient Calibration (GC) mechanism inspired by Early Learning Regularization (ELR) [21], which is originally designed for MCC, to alleviate overfitting to false negatives.
-To further improve generalizability, we incorporate the effective data augmentation technique Mixup [15] within the GC process.
-Finally, to ensure the effectiveness of this calibration process, we conduct an in-depth analysis of learning dynamics in SPML and propose a simple yet theoretically grounded indicator for adaptive activation of GC after an initial warm-up stage.
-This mechanism is crucial for avoiding the negative influence of low-quality pseudo-labels on model optimization.
-We evaluate the AdaGC on two benchmark RS multi-label datasets with two single-positive label simulation ways.
-Extensive experiments demonstrate the effectiveness and robustness of AdaGC.
+First, we propose a theoretically grounded Gradient Calibration (GC) mechanism designed specifically for SPML.
+GC aims to alleviate the model’s tendency to overfit false negatives by correcting gradients using pseudo-label guidance.
+Unlike Early Learning Regularization (ELR) [liu_early-learning_2020-2], which is formulated for multi-class classification under the simplex constraint, GC is rigorously extended to the SPML setting without such a constraint.
+Second, to guarantee pseudo-label quality, we introduce a dual exponential moving average (EMA) module for robust pseudo-label estimation, which leverages the consensus between the temporally smoothed predictions of the student model and the teacher model’s outputs.
+The former applies EMA to predictions, and the latter to model weights.
+Then, to ensure the effectiveness of GC, we conduct an in-depth analysis of learning dynamics in SPML and propose a theoretically grounded indicator for adaptive activation of Gradient Calibration (GC) after an initial warm-up stage.
+This mechanism is central to our design, which is capable of preventing interference from underfitting or noise-driven overfitting to stabilize optimization.
+We also incorporate the Mixup data augmentation technique [Zhang_2018_ICLR] within GC to further improve model’s generalizability.
+Extensive experiments on two benchmark RS multi-label datasets demonstrate the effectiveness and robustness of AdaGC.
 The main contributions of this paper are summarized as follows:
-  • We propose a two-stage SPML framework for RS imagery based on a student–teacher architecture, incorporating an early learning detection strategy that guides the activation of GC to combat overfitting to label noise.
-Furthermore, we provide theoretical analyses that support the effectiveness of our approach in SPML scenarios.
+  • We propose a two-stage RS SPML framework, which incorporates an early learning detection strategy to guide the activation of GC to combat underfitting and overfitting to label noise.
+The theoretical analyses are also provided to support the effectiveness of the framework design.
   • We introduce a dual-EMA pseudo-label generation strategy that enhances label completeness and reliability by leveraging temporal prediction fusion.
-  • We perform extensive empirical evaluations on two benchmark RS datasets, covering high- and low-resolution imagery, and two single positive label noise simulation strategies (Random and Dominant).
-We also provide a comprehensive benchmarking of existing SPML methods, offering valuable insights for future research in this domain.
+  • We conduct extensive experiments on two benchmark RS datasets covering both high- and low-resolution imagery.
+For low-resolution data, we employ Random and Dominant SPML noise simulations to better reflect real-world annotation behavior.
+For high-resolution imagery, where identifying a dominant class is inherently difficult, we rely on Random simulation supplemented with additional Manual single-positive annotations.
+We provide a comprehensive benchmarking of existing SPML methods, offering valuable insights for future research in this domain.
 The remainder of this paper is organized as follows.
@@
 II Related Works
-This section reviews advancements in MLC, first focusing on its application in RS and then discussing in general the challenges posed by noisy labels.
-Finally, we delve into SPML, detailing various methods developed in CV to address its unique difficulties.
+This section reviews advancements in MLC, with emphasis on RS applications, challenges from noisy labels, and recent SPML methods developed in CV.
 II-A Remote Sensing Image Multi-label Classification
-Early DL research in MLC for RS imagery primarily focused on high-resolution data and employed Convolutional Neural Networks for feature extraction, followed by classifiers such as Radial Basis Function Neural Networks [23] or Structured Support Vector Machines [24].
-These approaches often relied on conventional transfer learning, using CNNs pre-trained on generic CV datasets (e.g., ImageNet [25]) as fixed feature extractors.
+Early DL research in MLC for RS imagery primarily focused on high-resolution data and employed Convolutional Neural Networks for feature extraction, followed by classifiers such as Radial Basis Function Neural Networks [Zeggada_2017_GRSL] or Structured Support Vector Machines [Koda_2018_TGRS].
+These approaches often relied on conventional transfer learning, using CNNs pre-trained on generic CV datasets (e.g., ImageNet [Russakovsky_2015_IJCV]) as fixed feature extractors.
 However, the significant differences between natural CV scenes and complex RS imagery can lead to suboptimal classification performance, limiting the effectiveness of such direct transfer.
@@
 For instance, Stivaktakis et al.
-[26] introduced a data augmentation scheme to facilitate end-to-end training of shallow CNNs, along with replacing the softmax layer with a sigmoid function to accommodate multi-label outputs.
+[Stivaktakis_2019_GRSL] introduced a data augmentation scheme to facilitate end-to-end training of shallow CNNs, along with replacing the softmax layer with a sigmoid function to accommodate multi-label outputs.
 While this direct adaptation remains a common practice in multi-label scenarios, it may result in the imprecise identification of multiple classes.
@@
 Hua et al.
-[8] proposed a class-wise attention-based Recurrent Neural Network (RNN) to explicitly model the co-occurrence relationships among multiple classes, thus generating class predictions sequentially.
-Similarly, an attention-aware label relational reasoning network was introduced in [13] to localize discriminative regions and characterize inter-label relationships based on feature maps.
+[Hua_2019_JPRS] proposed a class-wise attention-based Recurrent Neural Network (RNN) to explicitly model the co-occurrence relationships among multiple classes, thus generating class predictions sequentially.
+Similarly, an attention-aware label relational reasoning network was introduced in [hua_relation_2020] to localize discriminative regions and characterize inter-label relationships based on feature maps.
 Alshehri et al.
-[27] presented an encoder-decoder neural network architecture, where the encoder includes a squeeze-and-excitation layer to model channel-wise dependencies, and the RNN-based decoder serves as an adaptive spatial attention module.
+[Alshehri_2019_ACCESS] presented an encoder-decoder neural network architecture, where the encoder includes a squeeze-and-excitation layer to model channel-wise dependencies, and the RNN-based decoder serves as an adaptive spatial attention module.
 These attention-driven strategies have demonstrated considerable success in identifying informative areas through attention maps derived from convolutional features.
@@
 A major hurdle in MLC in RS is the difficulty in obtaining exhaustive multi-label annotations, which often results in noisy labels.
-This noise can be categorized as either missing labels (subtractive noise), where a class present in the considered image is not annotated, or as wrong labels (additive noise), where an absent class is incorrectly annotated [10].
+This noise can be categorized as either missing labels (subtractive noise), where a class present in the considered image is not annotated, or as wrong labels (additive noise), where an absent class is incorrectly annotated [Burgert_2022_TGRS].
 Such multi-label noise, which can be introduced through zero-cost annotation (e.g., using thematic products), can severely impair model training.
-Consequently, developing methods robust to multi-label noise in RS is an active research area, usually categorized as weakly supervised multi-label learning (WSML) [28].
-Methods for handling noisy labels in multi-label settings often draw inspiration from related fields, such as positive-unlabeled learning (PUL) [29, 30, 31] and noisy multi-class learning [32].
-Common strategies involve developing robust loss functions that are less sensitive to label noise or class imbalance [33].
-Another approach is pseudo-labeling, where likely labels for unobserved or unknown annotations are inferred to provide additional training signals [34].
+Consequently, developing methods robust to multi-label noise in RS is an active research area, usually categorized as weakly supervised multi-label learning (WSML) [Xie_2023_TPAMI].
+Methods for handling noisy labels in multi-label settings often draw inspiration from related fields, such as positive-unlabeled learning (PUL) [duPlessis_2014_NIPS, Gang_2016_NIPS, bekker_2020_ML] and noisy multi-class learning [perantonirobust].
+Common strategies involve developing robust loss functions that are less sensitive to label noise or class imbalance [li2020robust].
+Another approach is pseudo-labeling, where likely labels for unobserved or unknown annotations are inferred to provide additional training signals [Zhang_2020_RS].
 The “memorization effect,” where models tend to learn clean labels early but later start to memorize noisy ones, has been widely observed in noisy MCC.
-ELR [21] was developed to specifically counteract this, by adjusting the influence of gradients from potentially noisy samples, particularly in the early stages of training.
+Early Learning Regularization (ELR) [Liu_2020_NIPS] was developed to specifically counteract this, by adjusting the influence of gradients from potentially noisy samples, particularly in the early stages of training.
 Burgert et al.
-[10] later extended it to MLC in a straightforward manner and empirically demonstrated its effectiveness.
-Additional strategies include sample selection, which aims to identify and potentially exclude samples with noisy labels [35], and label correction/refurbishment, which attempts to correct the noisy labels (e.g., using collaborative learning frameworks [36, 37]).
-Other methods, such as those in the large loss (LL) [16] family, including LL-R, LL-Ct, and LL-Cp, are developed to recognize and handle mislabeled samples with large losses, based on the observation that noisy labels often lead to high loss values.
+[Burgert_2022_TGRS] later extended it to MLC in a straightforward manner and empirically demonstrated its effectiveness.
+Additional strategies include sample selection, which aims to identify and potentially exclude noisy labels [Paris_2021_TGRS], label correction/refurbishment, which attempts to correct noisy labels (e.g., using collaborative learning frameworks [Aksoy_2021_ICIP, Aksoy_2024_TNNLS]), and noise-robust regularization, such as using word embedding similarities to regularize model training [hua_learning_2020].
+Other methods, such as those in the large loss (LL) [Kim_2022_CVPR] family, including LL-R, LL-Ct, and LL-Cp, are developed to recognize and handle mislabeled samples with large losses, based on the observation that noisy labels often lead to high loss values.
 Specifically, LL-R rejects samples with large losses, LL-Ct applies temporary corrections, and LL-Cp employs permanent corrections.
-BoostLU [19] complements LL by enhancing predicted positive regions via a function applied to class activation maps, further improving the robustness.
+BoostLU [Kim_2023_CVPR] complements LL by enhancing predicted positive regions via a function applied to class activation maps, further improving the robustness.
 In WSML for RS, subtractive noise, i.e., missing true labels, is a prevalent issue.
@@
 III Methodology
-Figure 2: Flowchart of the proposed Adaptive Gradient Calibration (AdaGC) method for single positive multi-label learning in remote sensing image classification.
+Figure 2: Flowchart of the proposed Adaptive Gradient Calibration (AdaGC) method for single-positive multi-label learning in remote sensing image classification.
+Pseudo-labels $\mathbf{t}$ are generated by combining the teacher model’s predictions $\mathbf{p}^{T}$ and the student model’s predictions $\tilde{\mathbf{p}}^{S}$ according to (20).
+EMA is also applied to the student model’s predictions during the warm-up stage, which is omitted for simplification.
 As illustrated in Fig.
-2, the proposed method adopts a student–teacher framework, where the teacher model’s weights are updated via EMA.
+2, the proposed method adopts a student–teacher architecture, where the teacher model’s weights are updated via EMA.
 The training pipeline consists of two stages.
 In the first stage, the student model is trained directly on noisy labels, during which the teacher model’s validation accuracy (with respect to noisy labels) is monitored as an early learning indicator to determine when to activate the second stage— Gradient Calibration (GC).
-Next, we first define the problem before presenting the technical details of the proposed method.
+Next, we first define the problem before presenting the technical details of AdaGC.
 III-A Problem Definition
@@
 III-B Gradient Calibration
-The concept of gradient calibration for handling noisy labels was first proposed in [21] for MCC, where ELR introduces an auxiliary regularization term to constrain the standard cross-entropy loss, thus preventing the memorization of noisy labels.
+The concept of gradient calibration for handling noisy labels was first proposed in [Liu_2020_NIPS] for MCC, where ELR introduces an auxiliary regularization term to constrain the standard cross-entropy loss, thus preventing the memorization of noisy labels.
 The regularization term for MCC is defined as follows:
-$\mathcal{R}_{\text{MCC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\log\left(1-\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle\right),$ (3)
-where $n$ is the number of training samples, $\mathbf{p}$ denotes the softmax output vector of the multi-class model, and $\mathbf{t}$ represents the soft pseudo-labels.
+$\mathcal{R}_{\text{MCC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\log\left(1-\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle\right).$ (3)
+Here, both $\mathbf{p}_{i}$ and $\mathbf{t}_{i}$ are probability vectors whose elements are non-negative and sum to 1.
+This simplex constraint ensures that their inner product $\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle$ lies within the range $[0,1]$.
 The gradient of (3) is defined as follows:
@@
 with $\mathbf{g}_{i}\in\mathbb{R}^{C}$, whose values are computed as follows:
-$\mathbf{g}_{i,c}=\frac{p_{i,c}}{1-\langle\mathbf{p}_{i},\mathbf{t}_{i}\rangle}\sum_{k=1}^{C}\left(t_{i,k}-t_{i,c}\right)\cdot p_{i,k}.$ (5)
+$g_{i,c}=\frac{p_{i,c}}{1-\langle\mathbf{p}_{i},\mathbf{t}_{i}\rangle}\sum_{k=1}^{C}\left(t_{i,k}-t_{i,c}\right)\cdot p_{i,k}.$ (5)
 As indicated by (5), the sign of (4) is governed by $t_{i,k}-t_{i,c}$.
@@
 To address this limitation, we emphasize the gradient calibration on the potential positive labels with the following regularization term:
-$\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\log\left(1-p_{i,c}\cdot t_{i,c}\right).$ (7)
+$\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\mathbb{I}(y_{i,c}=0)\log\left(1-p_{i,c}\cdot t_{i,c}\right),$ (7)
+where $\mathbb{I}(\texttt{True})=1$ otherwise 0.
+Considering the SPML setting, GC is applied only to negative labels, excluding its use for positive ones in practice.
 We can compute the gradient of (7) as follows:
-$\nabla_{\theta}\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\nabla_{\theta}f(\mathbf{x}_{i};\theta)\cdot g_{i,c},$ (8)
+$\nabla_{\theta}\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\mathbb{I}(y_{i,c}=0)\nabla_{\theta}f_{c}(\mathbf{x}_{i};\theta)\cdot g_{i,c},$ (8)
 where $g_{i,c}$ is defined as follows:
-$g_{i,c}=\frac{p_{i,c}-1}{\frac{1}{p_{i,c}\cdot t_{i,c}}-1}\leq 0.$ (9)
-We argue that (7) is particularly suitable for the SPML setting, as it specifically penalizes false negatives.
-Accordingly, we apply GC only to negative labels, excluding its use for positive ones in practice.
+$g_{i,c}=\frac{-t_{i,c}}{1-p_{i,c}\cdot t_{i,c}}\cdot{p_{i,c}(1-p_{i,c})}\leq 0,$ (9)
+where $p_{i,c}(1-p_{i,c})$ is the sigmoid derivative.
+We argue that (7) is particularly suitable for the SPML setting, as it specifically penalizes false negatives by pushing the model to predict positive scores for (negative) labels with higher pseudo-label scores $t_{i,c}$.
 The final objective is a combination of the binary cross-entropy loss and the GC regularization:
@@
 Liu et al.
-[22] show that adaptively triggering correction before overfitting to label noise is crucial for the effectiveness of sample-correction-related techniques.
+[liu_aio2_2024] show that adaptively triggering correction before overfitting to label noise is crucial for the effectiveness of sample-correction-related techniques.
 Inspired by this, in Section III-C we introduce a simple early learning detection mechanism to appropriately activate GC.
@@
 III-C Early Learning Detection
-Pseudo-labelling methods typically require an initial warm-up phase, whose duration, often determined empirically, critically impacts performance.
-In SPML, premature pseudo-label initiation introduces additive noise, whereas excessive delay risks overfitting to inherent subtractive label noise, thus limiting false negative detection.
-Optimal pseudo-labelling activation is crucial for model performance.
-Monitoring performance metrics on a clean validation set is ideal, as model performance typically improves, plateaus, then declines (indicating overfitting) [40, 41].
-The plateau, detectable by the slope of performance metrics, identifies the optimal time to activate pseudo-labelling [22].
-However, SPML generally lacks clean validation data, typically offering only single positive labels, akin to the training set.
-While standard metrics on noisy sets might differ, we demonstrate that the mean average precision (mAP) computed on the noisy validation set mirrors the behavior of mAP* computed on the clean validation set, showing approximately consistent trends throughout training.
-The mAP is the macro average of average precision (AP) for each class:
+Pseudo-labelling methods typically require an initial warm-up phase to mitigate the trade-off between additive noise (premature initiation) and subtractive noise overfitting (excessive delay).
+Ideally, the optimal activation time is identified by the performance plateau on a clean validation set [Arpit_2017_ICML, Zhang_2021_CommACM, liu_aio2_2024].
+In SPML, where clean validation data is unavailable, we propose monitoring the mean average precision (mAP) on the noisy validation set.
+Below, we provide a formal justification for this criterion by analyzing the bounds and stationarity of the noisy metric.
+III-C1 Theoretical Framework
+(a) Random SPML
+(b) Dominant SPML
+Figure 3: True and noisy validation mAP trends of the teacher model during training on refined BigEarthNet [clasen_reben_2025_arxivonly]: (a) random SPML, (b) dominant SPML.
+The trends align with Propositions 1 and 2, and confirm the validity of the stationarity theorem.
+Let $\text{Rec}^{*}_{c}$ and $\text{Prec}^{*}_{c}$ denote the clean recall and precision for class $c$.
+In SPML, the set of real positives $\text{P}_{c}$ is reduced by the flipped positives $\text{F}_{c}=\beta_{c}\text{P}_{c}$, where $\beta_{c}\in(0,1)$ is the noise rate.
+Defining $\text{PF}_{c}=\alpha_{c}\cdot\text{F}_{c}$, where $\alpha_{c}\in[0,1]$, as the proportion of correctly classified flipped positives, the noisy metrics relate to the clean ones as follows:
+$\begin{split}\text{Rec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{P}_{c}-\text{F}_{c}}=\frac{\text{Rec}^{*}_{c}-\alpha_{c}\beta_{c}}{1-\beta_{c}},\\ \text{Prec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{PP}_{c}}=\text{Prec}^{*}_{c}\left(1-\frac{\alpha_{c}\beta_{c}}{\text{Rec}^{*}_{c}}\right),\end{split}$ (11)
+where $\text{TP}_{c}$ and $\text{PP}_{c}$ are the true and predicted positives, respectively. The mAP is the macro average of average precision (AP) for each class:
 $\text{mAP}=\frac{1}{C}\sum_{c=1}^{C}{\text{AP}_{c}},$ (12)
 where $\text{AP}_{c}=\int_{0}^{1}{\text{Prec}_{c}\left(\text{Rec}_{c}\right)d\text{Rec}_{c}}$.
-For a clean validation set, recall $\text{Rec}^{*}_{c}$ and precision $\text{Prec}^{*}_{c}$ are defined as follows:
-$\begin{split}\text{Rec}^{*}_{c}=\frac{\text{TP}_{c}}{\text{P}_{c}}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FN}_{c}},\\ \text{Prec}^{*}_{c}=\frac{\text{TP}_{c}}{\text{PP}_{c}}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}},\end{split}$ (12)
-where $\text{TP}_{c}$, $\text{FP}_{c}$, and $\text{FN}_{c}$ are the true positives, false positives, and false negatives for class $c$, respectively, while $\text{P}_{c}=\text{TP}_{c}+\text{FN}_{c}$ and $\text{PP}_{c}=\text{TP}_{c}+\text{FP}_{c}$ are the real and predicted positives.
-In SPML, flipped positives are $\text{F}_{c}=\beta_{c}\cdot\text{P}_{c}$, where $\beta_{c}\in(0,1)$ is the proportion of flipped positives.
-Let $\text{PF}_{c}=\alpha_{c}\cdot\text{F}_{c}$ be the proportion of correctly classified flipped positives, where $\alpha_{c}\in[0,1]$.
-Thus, noisy validation metrics $\text{Rec}_{c}$ and $\text{Prec}_{c}$ are functions of their clean counterparts:
-$\begin{split}\text{Rec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{P}_{c}-\text{F}_{c}}=\frac{\text{Rec}^{*}_{c}-\alpha_{c}\cdot\beta_{c}}{1-\beta_{c}},\\ \text{Prec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{PP}_{c}}=\text{Prec}^{*}_{c}\cdot\left(1-\frac{\alpha_{c}\cdot\beta_{c}}{\text{Rec}^{*}_{c}}\right).\end{split}$ (13)
-These metrics depend on the label noise flip rate $\beta_{c}$ and the classifier’s behavior on flipped labels $\alpha_{c}$, which can vary, altering the noisy validation dynamic from the clean one.
-However, assumptions about $\alpha_{c}$ based on noise type allow us to draw some conclusions.
-We consider two SPML noise types: Random and Dominant.
-In both cases, we will demonstrate below that the expected noisy $\text{AP}_{c}$ for class $c$ relates to its true counterpart $\text{AP}_{c}^{*}$ via a class-specific scaling factor $\omega_{c}$:
-$\mathbb{E}[\text{AP}_{c}]=\omega_{c}\cdot\text{AP}_{c}^{*}.$ (14)
-Consequently, the expected noisy mAP, can be expressed relative to the true mAP* as follows:
-$\mathbb{E}[\text{mAP}]=\frac{1}{C}\sum_{c=1}^{C}{\omega_{c}\cdot\text{AP}_{c}^{*}}=\bar{\omega}\cdot\text{mAP}^{*}+\text{Cov}\left(\omega,\text{AP}^{*}\right),$ (15)
-where $\bar{\omega}=\frac{1}{C}\sum_{c=1}^{C}{\omega_{c}}$ is the mean scaling factor across classes, and $\text{Cov}\left(\omega,\text{AP}^{*}\right)$ denotes the covariance between the class-specific scaling factors and the true APs.
-The sign and magnitude of this covariance term determine the deviation from a simple mean-scaling relationship.
-III-C1 Random SPML
-In random SPML, single positive labels are chosen uniformly at random from the set of true positives for each instance.
-This type of noise is generally considered instance-independent [42].
-While the trained model may inevitably fit this noise in the training labels, it cannot meaningfully generalize to it or predict the noisy labels in the validation set.
-Under this assumption, the proportion of correctly classified flipped positives, $\alpha_{c}$, on average equals the recall for that class, i.e., $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$, from which we can derive the following noisy validation metrics:
-$\mathbb{E}\left[\text{Rec}_{c}\right]\approx\text{Rec}^{*}_{c},\quad\mathbb{E}\left[\text{Prec}_{c}\right]\approx\left(1-\beta_{c}\right)\cdot\text{Prec}^{*}.$ (16)
-Substituting these into the AP definition, we find that the expected noisy AP for each class is scaled down by the non-flip rate:
-$\mathbb{E}[\text{AP}_{c}]\approx(1-\beta_{c})\cdot\text{AP}_{c}^{*}.$ (17)
-Here, the scaling factor is $\omega_{c}=(1-\beta_{c})$. Substituting this into the general mAP relationship in (15), we get the following:
-$\begin{split}\mathbb{E}[\text{mAP}]&\approx\frac{1}{C}\sum_{c=1}^{C}{(1-\beta_{c}})\cdot\text{AP}_{c}^{*}<\text{mAP}^{*},\\ \mathbb{E}[\text{mAP}]&\approx(1-\bar{\beta})\cdot\text{mAP}^{*}-\text{Cov}\left(\beta,\text{AP}^{*}\right),\end{split}$ (18)
-where $\bar{\beta}$ is the mean flip rate across classes.
-This implies that the expected noisy mAP will be lower than the true mAP*.
-The covariance term $\text{Cov}\left(\beta,\text{AP}^{*}\right)$ captures how the class-specific flip rates correlate with their true performances.
-In general, we can expect that better-performing classes experience less flipping, implying negative covariance and therefore a mitigation effect on the mean-scaling relationship in (18), i.e., $(1-\bar{\beta})\cdot\text{mAP}^{*}\leq\mathbb{E}[\text{mAP}]<\text{mAP}^{*}$.
-III-C2 Dominant SPML
-In contrast to random SPML, dominant SPML belongs to the category of instance-dependent label noise [42].
-Here, the single-positive label is consistently the “dominant” one among the true positives.
-This allows the model to potentially learn this label-noise model, which can be reflected in the noisy validation mAP later in training.
-This results in three consecutive trends during learning [22]: i) an early learning phase where $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$ still holds, yielding similar results to random SPML; ii) a transition phase; and iii) a label-noise model fitting phase.
-Once the model learns the label-noise behavior, the expected proportion of correctly classified flipped positives approaches zero, as the model learns to ignore non-dominant true positives, i.e., $\mathbb{E}\left[\alpha_{c}\right]\to 0$.
-This implies the following expected noisy validation metrics:
-$\mathbb{E}[\text{Rec}_{c}]\approx\frac{\text{Rec}_{c}^{*}}{1-\beta_{c}},\quad\mathbb{E}[\text{Prec}_{c}]\approx\text{Prec}_{c}^{*},$ (19)
-which leads to the following expected AP relationship:
-$\mathbb{E}[\text{AP}_{c}]\approx\frac{1}{1-\beta_{c}}\cdot\text{AP}_{c}^{*}.$ (20)
-In this case, the scaling factor is $\omega_{c}={1}/(1-\beta_{c})$. The corresponding mAP relationship is as follows:
-$\small\begin{split}\text{mAP}^{*}&<\mathbb{E}[\text{mAP}]=\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}\cdot\text{AP}_{c}^{*}}<1,\\ \mathbb{E}[\text{mAP}]&\approx\left(\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}}\right)\text{mAP}^{*}+\text{Cov}\left(\frac{1}{1-\beta},\text{AP}^{*}\right).\end{split}$ (21)
-This indicates that the expected noisy mAP will be higher than the true $\text{mAP}_{c}^{*}$.
-The covariance term captures if classes with higher true APs are affected by the scaling.
-In general, we expect highly dominant classes to perform better than others.
-This leads to a negative covariance value, i.e., $\text{mAP}^{*}<\mathbb{E}[\text{mAP}]\leq\left(\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}}\right)\text{mAP}^{*}$.
-Consequently, noisy $\text{AP}_{c}$ are initially lower than $\text{AP}_{c}^{*}$ but later they increase their value.
-The noisy $\text{AP}_{c}$ approximately remain proportional to $\text{AP}_{c}^{*}$ during these phases, with the “transition phase” marking a departure from this simple scaling, possibly shifting the plateau’s appearance between them.
-III-C3 Robust Plateau Detection using Teacher Model
-(a) Random SPML
-(b) Dominant SPML
-Figure 3: True and noisy validation mAP trends of the teacher model during the training with AN on the refined BigEarthNet [14] dataset: (a) random SPML, and (b) dominant SPML.
-As demonstrated in (18) and (21), although a proportional relationship generally holds between the noisy mAP and the true $\text{mAP}^{*}$, the bias term (i.e., the covariance) can introduce deviations from this relationship.
-However, we argue that this bias term rarely compromises the analysis, especially in the early learning stage.
-At the beginning, the model primarily fits correctly labelled samples [21], leading to a weak correlation between $\beta_{c}$ and $\text{AP}_{c}^{*}$.
-As a result, the scaling relationship predominantly governs (18) and (21).
-Later, when the $\text{AP}_{c}^{*}$ scores plateau, the covariance term’s contribution to the slope of the noisy mAP becomes negligible, as it exhibits little to no variation.
-Together, these factors enable the noisy mAP to closely mirror the true performance plateau, making it a reliable indicator for identifying the end of the early learning phase.
-We propose monitoring the teacher model’s epoch-wise noisy validation mAP score to ensure a stable and reliable detection.
-The teacher model is maintained as an EMA of the student model’s weights.
-It acts as a low-pass filter, smoothing the student’s unstable learning dynamics with fluctuations caused by noisy gradients or erratic updates.
-As shown in Fig.
-3, the noisy and true mAP scores computed from the teacher model closely follow similar plateau trends, consistent with the relationships established in (18) and (21).
-This smoothed behavior allows for a more accurate and robust detection of the optimal epoch for pseudo-labeling.
-In practice, we online select the epoch that achieves the highest noisy validation mAP score from the teacher model, provided there is no significant improvement within the following $b_{e}$ epochs.
-Once this optimal epoch is detected, the model weights are automatically restored from that epoch, and pseudo-labeling is initiated concurrently with GC.
+Proposition 1 (Random SPML Lower Bound). Under instance-independent noise, the expected noisy mAP is a lower bound of the clean $\text{mAP}^{*}$.
+Proof: In Random SPML, the noise is independent of the input features [Perantoni_2022_TGRS].
+Thus, the model cannot generalize to the random flips, implying $\mathbb{E}[\alpha_{c}]\approx\text{Rec}^{*}_{c}$.
+Substituting this into (11) yields $\mathbb{E}[\text{AP}_{c}]\approx(1-\beta_{c})\text{AP}_{c}^{*}$.
+The expected noisy mAP becomes:
+$\mathbb{E}[\text{mAP}]\approx(1-\bar{\beta})\text{mAP}^{*}-\text{Cov}(\beta,\text{AP}^{*}),$ (13)
+where $\bar{\beta}$ is the mean flip rate. Since better-performing classes generally experience less flipping (negative covariance), $\mathbb{E}[\text{mAP}]<\text{mAP}^{*}$. $\square$
+Fig. 3(a) shows the trends observed while training on the refined BigEarthNet [clasen_reben_2025_arxivonly] dataset with random single-positive labels, which are aligned with our theoretical analysis.
+Proposition 2 (Dominant SPML Upper Bound). Under instance-dependent dominant noise, the expected noisy mAP is an upper bound of the clean $\text{mAP}^{*}$.
+Proof: In Dominant SPML, the model learns the label-noise model [Perantoni_2022_TGRS], ignoring non-dominant true positives ($\mathbb{E}[\alpha_{c}]\to 0$).
+Eq.
+(11) simplifies to $\mathbb{E}[\text{Rec}_{c}]\approx\frac{\text{Rec}_{c}^{*}}{1-\beta_{c}}$ and $\mathbb{E}[\text{Prec}_{c}]\approx\text{Prec}_{c}^{*}$, yielding $\mathbb{E}[\text{AP}_{c}]\approx\frac{1}{1-\beta_{c}}\text{AP}_{c}^{*}$.
+The noisy mAP becomes:
+$\mathbb{E}[\text{mAP}]\approx\overline{\left(\frac{1}{1-\beta}\right)}\text{mAP}^{*}+\text{Cov}\left(\frac{1}{1-\beta},\text{AP}^{*}\right).$ (14)
+Assuming dominant classes perform better (negative covariance), $\text{mAP}^{*}<\mathbb{E}[\text{mAP}]$. $\square$
+Generally, Dominant SPML results in three consecutive trends during learning [liu_aio2_2024]: i) an early learning phase where $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$ still holds, yielding similar results to random SPML; ii) a transition phase; and iii) a label-noise model fitting phase.
+Fig.
+3(b) plots the trends observed while training on the refined BigEarthNet [clasen_reben_2025_arxivonly] dataset with dominant single-positive labels, showing this three-phase phenomenon.
+III-C2 Adaptive Triggering Mechanism
+We now address the validity of using the noisy metric for early stopping.
+Theorem 1 (Stationarity Preservation). The stationary points of the expected noisy mAP coincide with those of the clean $\text{mAP}^{*}$ in the late early-learning phase.
+Proof: Let $t$ denote the training epoch.
+Combining Propositions 1 and 2, the noisy metric is a linear transformation of the clean metric: $\mathbb{E}[\text{mAP}(t)]=\bar{\omega}\text{mAP}^{*}(t)+\text{Cov}(\omega,\text{AP}^{*}(t))$.
+Taking the temporal derivative:
+$\frac{\partial\mathbb{E}[\text{mAP}]}{\partial t}=\bar{\omega}\frac{\partial\text{mAP}^{*}}{\partial t}+\frac{\partial}{\partial t}\text{Cov}(\omega,\text{AP}^{*}(t)).$ (15)
+The noise parameters $\omega$ are constant.
+Near the performance plateau ($t\to t^{*}$), class-wise learning dynamics tend to stabilize, making the covariance term approximately constant at that time ($\frac{\partial}{\partial t}\text{Cov}\approx 0$) Consequently:
+$\frac{\partial\mathbb{E}[\text{mAP}]}{\partial t}\propto\frac{\partial\text{mAP}^{*}}{\partial t}.$ (16)
+Thus, identifying the peak of the noisy mAP is mathematically equivalent to identifying the peak of the clean mAP. $\square$
+Implementation: To mitigate stochastic fluctuations in the single-run estimates of $\mathbb{E}[\text{mAP}]$, we monitor the mAP of the teacher model (maintained via EMA), which acts as a low-pass filter on the model’s weights (see Fig.
+6).
+We select the epoch maximizing the teacher’s noisy validation mAP (with a patience of $b_{e}$ epochs) to initiate pseudo-labeling.
 III-D Dual-EMA Pseudo-Label Generation
 As discussed in Section III-B, GC relies on high-quality pseudo-labels.
-Since it is triggered after early learning detection on the teacher model, the teacher’s predictions are a natural choice.
+The teacher’s predictions are a natural choice, since the early learning detection is on the teacher model.
 However, due to slow weight updates, the teacher model may exhibit a delay in incorporating recently acquired knowledge after GC begins.
 Therefore, we propose to combine predictions from both teacher and student models.
-Specifically, we apply EMA to the student’s predictions, using a smaller smoothing factor than that for the teacher model, to avoid overfitting to its own outputs while maintaining adaptability.
+Specifically, we apply EMA to the student’s predictions with a smaller smoothing factor than that for the teacher model, to avoid overfitting to its own outputs while maintaining adaptability.
 Let $\mathbf{p}_{i}^{T}$ and $\mathbf{p}_{i}^{S}$ denote the prediction probabilities of the teacher and student models for sample $i$, respectively. For the teacher model, we have the following:
@@
 $\theta^{T}(e)=\beta_{t}\cdot\theta^{T}(e-1)+(1-\beta_{t})\cdot\theta^{S}(e)$ (18)
-with $\theta^{T}(e)$ and $\theta^{S}(e)$ being the model weights of the teacher and student models at epoch $e$, respectively.
-We set the smoothing coefficient $\beta_{t}=0.999$ following [43].
-Then, we compute the temporarily EMA-smoothed student predictions at the $e$th iteration with the coefficient $\beta_{s}\in[0,1)$ as follows:
+with $\theta^{T}(e)$ and $\theta^{S}(e)$ being the model weights of the teacher and student models at iteration $e$.
+We set the smoothing coefficient $\beta_{t}=0.999$ following [tarvainen_mean_2017].
+Then, the temporarily EMA-smoothed student predictions are as follows:
 $\tilde{\mathbf{p}}_{i}^{S}(e)=\beta_{s}\cdot\tilde{\mathbf{p}}_{i}^{S}(e-1)+(1-\beta_{s})\cdot\mathbf{p}_{i}^{S}(e).$ (19)
-We empirically set $\beta_{s}=0.8$ to guarantee the update speed of the student model’s prediction.
-Then, the pseudo label $\mathbf{t}_{i}(e)$ used for gradient calibration is computed as the combination of the teacher and student predictions as follows:
+We empirically set $\beta_{s}=0.8$ to guarantee the update speed of the student’s predictions.
+Then, the pseudo-label $\mathbf{t}_{i}(e)$ used for gradient calibration is computed as the combination of the teacher and student predictions as follows:
 $\mathbf{t}_{i}=\gamma\cdot\mathbf{p}_{i}^{T}+(1-\gamma)\cdot\tilde{\mathbf{p}}_{i}^{S},$ (20)
@@
 III-E Mixup
-To further enhance robustness against label noise and improve generalization, we incorporate Mixup [15] into the GC stage.
+To further enhance robustness against label noise and improve generalization, we incorporate Mixup [Zhang_2018_ICLR] into the GC stage.
 For each batch, we randomly sample a pairing input $(\mathbf{x}_{j},\mathbf{y}_{j},\mathbf{t}_{j})$ for every original input $(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{t}_{i})$, and compute:
@@
 This data augmentation encourages linear behavior between examples, mitigates overfitting, and stabilizes training under noisy labels.
-IV Experiments
-This section details the experimental setup employed to evaluate the proposed AdaGC approach against a comprehensive set of state-of-the-art (SOTA) SPML methods.
-We first describe the datasets utilized and the related specific SPML noise types considered.
-Subsequently, we outline the experimental setup, including hyperparameter setting, evaluation metrics, model architectures, and comparison methods considered in this study.
-Finally, we present the comparison results and an in-depth ablation study of AdaGC.
-IV-A Datasets
-We evaluate the proposed method on two multi-label RS benchmark datasets characterized by different modalities and spatial resolutions: refined BigEarthNet (reBEN) [44, 45, 14], derived from SAR and multispectral Sentinel-1/2 imagery with a spatial resolution of 10m, and AID-multilabel [13], collected from Google Earth aerial RGB imagery with resolutions ranging from 0.5m to 8m per pixel.
-reBEN focuses on land cover classification, while AID-multilabel includes both land cover categories and object-level semantic information.
-Due to the availability of CORINE Land Cover (CLC) [46] pixel-wise masks for reBEN, we were able to simulate the two types of label noise described in Section III-C.
-In contrast, AID’s high spatial resolution introduces greater semantic complexity, making it challenging to determine a dominant class.
-Therefore, we simulate only Random single-positive label noise on AID.
-IV-A1 Refined BigEarthNet
-It contains 549,488 Sentinel-1/2 image pairs of size of $120\times 120$ pixels covering 10 European countries, each aligned with a pixel-level CLC reference map.
-Multi-label annotations are derived from these maps, resulting in 19 land cover classes (see Table I).
-Following the splits in [14], we use a 2:1:1 ratio for training, validation, and testing.
-Each input has 14 channels by concatenating 2 Sentinel-1 and 12 Sentinel-2 L2A bands.
-Leveraging the reference maps, on this dataset we simulate both types of SPML label noise (Random and Dominant).
-Dataset statistics for the two cases are provided in Table I.
+III-F Complexity Analysis
+To assess practical feasibility, we analyze the computational complexity of AdaGC relative to a standard classification baseline.
+First, one can note that AdaGC is a training-only framework, i.e., the dual-EMA module is discarded after training.
+Consequently, the inference complexity (parameters and FLOPs) is identical to the student backbone, ensuring no additional latency or memory overhead during deployment.
+Regarding training efficiency, a standard iteration incurs a cost of approximately $3\times$ the forward pass FLOPs ($1\text{F}$ forward and $\sim 2\text{F}$ backward).
+Instead, AdaGC processes two views during the GC stage: the original batch and a Mixup-augmented batch.
+However, the original batch is utilized exclusively for generating robust pseudo-labels via the student and teacher networks, which requires only forward passes, while the model optimization relies on the Mixup batch.
+Therefore, the computationally intensive backward pass is strictly necessary only for this view.
+This design results in a total theoretical cost of $\sim 5\text{F}$ per iteration, representing a moderate $\sim 66\%$ overhead compared to the baseline ($3\text{F}$) exclusively during the GC stage.
+Note that Mixup is simply a linear combination of two images.
+Thus, its additional FLOPs are mathematically negligible compared to the rest of the pipeline.
+In terms of memory complexity, the overhead is modest.
+The teacher EMA maintains only a second copy of the backbone parameters (not activations), contributing less than 10% of total memory in typical large-scale models.
+The temporally smoothed student predictions are stored as image-wise labels, which also require minimal memory.
+Taken together, both FLOPs and memory considerations indicate that AdaGC is computationally feasible for large-scale remote sensing applications.
 TABLE I: Refined BigEarthNet Land Cover Classes and Related Flip
 Rates $\beta_{c}$ on the Training Subset in the Two Cases of
-Random and Dominant SPML Label Noise. Macro
-and Micro Averages Are Provided, as Well as
-the Average Number of True Positive
-Labels per Image
+Random and Dominant SPML Label Noise
 Class Name Random $\beta_{c}$ Dominant $\beta_{c}$ Support
-Urban fabric 0.73 0.84 33626
-Industrial or commercial units 0.72 0.91 6398
-Arable land 0.63 0.51 88236
-Permanent crops 0.70 0.74 15190
-Pastures 0.66 0.62 42968
-Complex cultivation patterns 0.73 0.77 51034
-Land principally occupied by agriculture, with significant areas of natural vegetation 0.73 0.86 60127
-Agro-forestry areas 0.64 0.54 15082
-Broad-leaved forest 0.69 0.65 64164
-Coniferous forest 0.66 0.54 76344
-Mixed forest 0.70 0.67 78791
-Natural grassland and sparsely vegetated areas 0.69 0.82 6984
-Moors, heathland and sclerophyllous vegetation 0.70 0.70 6329
-Transitional woodland, shrub 0.70 0.85 64198
+Urban fabric 0.73 0.84 33,626
+Industrial or commercial units 0.72 0.91 6,398
+Arable land 0.63 0.51 88,236
+Permanent crops 0.70 0.74 15,190
+Pastures 0.66 0.62 42,968
+Complex cultivation patterns 0.73 0.77 51,034
+Land principally occupied by agriculture, with significant areas of natural vegetation 0.73 0.86 60,127
+Agro-forestry areas 0.64 0.54 15,082
+Broad-leaved forest 0.69 0.65 64,164
+Coniferous forest 0.66 0.54 76,344
+Mixed forest 0.70 0.67 78,791
+Natural grassland and sparsely vegetated areas 0.69 0.82 6,984
+Moors, heathland and sclerophyllous vegetation 0.70 0.70 6,329
+Transitional woodland, shrub 0.70 0.85 64,198
 Beaches, dunes, sands 0.72 0.94 738
-Inland wetlands 0.67 0.78 11344
+Inland wetlands 0.67 0.78 11,344
 Coastal wetlands 0.70 0.70 670
-Inland waters 0.67 0.73 30466
-Marine waters 0.08 0.05 35914
+Inland waters 0.67 0.73 30,466
+Marine waters 0.08 0.05 35,914
 Micro average 0.65 0.65
 Macro average $\bar{\beta}$ 0.66 0.70
-Average #labels/image 2.9
+Average number of true positive labels per image 2.9
+IV Experiments
+This section presents the experimental setups, comparative results, and a comprehensive ablation study of AdaGC.
+IV-A Datasets
+We evaluate the proposed method on two multi-label RS benchmark datasets with different modalities and resolutions: refined BigEarthNet (reBEN) [Sumbul_ben_2019_IGARSS, sumbul_BigEarthNet-mm_2021, clasen_reben_2025_arxivonly], derived from SAR and multispectral Sentinel-1/2 imagery at 10m resolution, and AID-multilabel [hua_relation_2020], derived from high-resolution RGB images at 0.5–8m resolution.
+While reBEN focuses on land cover classification, AID-multilabel includes both land cover categories and object-level semantic information.
+IV-A1 reBEN
+TABLE II: AID Multilabel Classes and Related Flip Rates $\beta_{c}$ on
+the Training Subset in the Two Cases of Random
+and Manual SPML Label Noise
+Class Name Random $\beta_{c}$ Manual $\beta_{c}$ Support
+Airplane 0.83 0.80 59
+Bare-soil 0.80 0.78 882
+Buildings 0.82 0.81 1,316
+Cars 0.85 0.85 1,210
+Chaparral 0.76 0.85 79
+Court 0.86 0.78 203
+Dock 0.94 0.83 166
+Field 0.66 0.57 127
+Grass 0.78 0.80 1,368
+Pavement 0.83 0.84 1,367
+Sand 0.41 0.48 153
+Sea 0.77 0.67 133
+Ship 0.90 0.89 181
+Tanks 0.82 0.73 66
+Trees 0.79 0.80 1,467
+Water 0.83 0.84 510
+Micro average 0.81 0.81
+Macro average $\bar{\beta}$ 0.79 0.77
+Average number of true positive labels per image 5.2
+It contains 549,488 Sentinel-1/2 image pairs of size of $120\times 120$ pixels covering 10 European countries, each aligned with a pixel-level CORINE Land Cover (CLC) reference map.
+Multi-label annotations are derived from these maps, resulting in 19 land cover classes (see Table I).
+Following the splits in [clasen_reben_2025_arxivonly], we use a 2:1:1 ratio for training, validation, and testing.
+Each input has 14 channels by concatenating 2 Sentinel-1 and 12 Sentinel-2 L2A bands.
+Leveraging the reference maps, on this dataset we simulate both types of SPML label noise described in Section III-C (Random and Dominant).
+The Dominant case better reflects real-world single-positive annotations, as low-resolution images contain limited details, where the prevailing class is more likely to be observed while minor classes are ignored.
+Dataset statistics for the two cases are provided in Table I.
 IV-A2 AID-multilabel
-It extends the AID scene classification dataset [47] with multi-label annotations.
+It extends the AID scene classification dataset [xia_aid_2017] with multi-label annotations.
 The dataset comprises 3,000 high-resolution RGB aerial image patches ($600\times 600$ pixels) across 17 classes, collected from China, U.S., and several European countries.
 We exclude the extremely rare “mobile home” class (only 2 samples), resulting in 16 classes.
-Following [13], 20% of the data is held out for testing.
+Following [hua_relation_2020], 20% of the data is held out for testing.
 We then use 25% of the remaining training set for validation.
+Unlike reBEN, AID’s high spatial resolution allows multiple objects to be easily recognized, making it difficult to identify a single dominant class.
+Therefore, in addition to Random simulation, we manually annotated the AID dataset in a single-positive manner to assess model performance under realistic SPML conditions.
+For each image, four randomly shuffled candidate classes were displayed at a time to speed up annotation, with the option to refresh if none were suitable.
+Consecutive images were also prevented from being assigned to the same class to reduce potential annotation bias.
+This procedure makes the Manual labels behave similarly to random SPML labels, yet retain a slight human selection bias.
+The entire annotation process took approximately 4.5 hours.
+We also found and corrected several label inconsistencies in the original AID-multilabel data, most of which are from the meadow folder.
+The corrected multi-label annotations, the manual single-positive labels, and the annotation tool will be released publicly along with the AdaGC code.
 Dataset statistics are provided in Table II.
 For simplicity, in the following we use AID to refer to AID-multilabel.
-TABLE II: AID Multilabel Classes and Related Flip Rates $\beta_{c}$ on the
-Training Subset in the Case of Random SPML Label Noise.
-Macro and Micro Averages Are Also Provided, as Well
-as the Average Number of True Positive
-Labels per Image
-Class name Random $\beta_{c}$ Support
-Airplane 0.83 59
-Bare-soil 0.80 883
-Buildings 0.82 1313
-Cars 0.84 1211
-Chaparral 0.72 58
-Court 0.85 207
-Dock 0.91 163
-Field 0.74 129
-Grass 0.79 1366
-Pavement 0.81 1408
-Sand 0.47 153
-Sea 0.77 133
-Ship 0.86 176
-Tanks 0.83 65
-Trees 0.78 1441
-Water 0.84 509
-Micro average 0.81
-Macro average $\bar{\beta}$ 0.79
-Average #labels/image 5.2
-IV-B Experimental Setup
-We adopt ResNet-34 and ResNet-50 as backbone networks for the reBEN and AID datasets, respectively, with batch sizes of 512 and 32.
-We choose ResNet-34 for reBEN as it has fewer parameters and lower computational complexity, which significantly reduces the overall runtime of our experiments.
-Despite being lighter, it shares a similar architecture with ResNet-50 and generally achieves comparable performance.
-Moreover, since our primary goal is to evaluate the effectiveness of the proposed method relative to baseline settings, rather than to maximize absolute accuracy, ResNet-34 provides a suitable and efficient compromise.
-Given the AID dataset’s limited size and RGB data, we initialize the network with ImageNet-pretrained weights.
-In this setup, we consider the realistic scenario of absence of clean labels for training and hyperparameter tuning.
-Hence, all methods’ hyperparameters were tuned via grid search by evaluating their mAP score on the noisy validation set.
-To manage the computational burden associated with the large number of methods and hyperparameters, we performed this tuning process using a subset of the respective datasets (20% of training/validation sets for reBEN and 50% for AID) and for a reduced number of epochs (10 epochs instead of 70 as in the final runs).
-For AdaGC, we set the Mixup Beta distribution parameter $\alpha$ to 1, and the GC loss weight $\lambda$ to 3, 5, and 8 for reBEN-Random, reBEN-Dominant, and AID, respectively, selected via hyperparameters tuning.
-For the final evaluation, we utilized the optimal hyperparameters identified during the tuning phase.
-Each model was then trained on the full training dataset for a total of 70 epochs, after which each model was evaluated on the test dataset.
-For the proposed AdaGC approach, the noisy validation mAP from the teacher model, as discussed in Section III-C, was used to detect the optimal epoch for initiating GC with pseudo-label generation.
-To ensure the robustness and statistical reliability of our results, each method underwent three Monte Carlo runs, allowing us to quantify the variability in performance due to random weight initialization and dataset shuffling.
-We report the average values of all metrics along with their standard deviations (computed using an unbiased estimator).
-For comprehensive evaluation, in addition to mAP, we report two ranking-based metrics, i.e., coverage and ranking loss (Rankloss), and four accuracy-based metrics, i.e., overall accuracy (OA), mean F1 score (mF1), mean precision (mprecision), and mean recall (mrecall).
-The metrics are defined as follows.
-  • Coverage measures how far we need to go down the ranked label list to cover all true labels:
-    $\frac{1}{N}\sum_{i=1}^{N}\max_{y\in\mathbf{y}_{i}}\text{rank}_{f}(\mathbf{x}_{i},y)-1,$ (27)
-    where $\text{rank}_{f}(x_{i},y)$ is the position of the ground-truth (GT) label $y_{i,c}$ in the ranking list predicted by $f$;
-  • Ranking loss computes the average fraction of label pairs that are incorrectly ordered:
-    $\small\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathbf{y}_{i}||\overline{\mathbf{y}_{i}}|}\sum_{(y_{1},y_{0})\in\mathbf{y}_{i}\times\overline{\mathbf{y}_{i}}}\mathbb{I}(f(\mathbf{x}_{i},y_{1})\leq f(\mathbf{x}_{i},y_{0})),$ (28)
-    where $\mathbf{y}_{i}$ and $\overline{\mathbf{y}_{i}}$ are the sets of relevant (positive) and irrelevant (negative) labels, respectively, and $\mathbb{I}(\texttt{True})=1$ otherwise 0;
-  • OA measures the percentage of samples with all labels correctly predicted:
-    $\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}\mathbb{I}(\hat{\mathbf{y}}_{i,c}=\mathbf{y}_{i,c});$ (29)
-  • mF1 is the average of F1-scores computed per class:
-    $\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Prec}_{c}\cdot\text{Rec}_{c}}{\text{Prec}_{c}+\text{Rec}_{c}}.$ (30)
-We compared AdaGC against a wide array of existing methods, categorized as follows.
-  • Baseline methods: We considered several baselines [12], such as AN, AN-LS, Weak (WAN), and EPR.
-All use a BCE loss at their core, while EPR uses only the BCE positive term and adds a regularization term, assuming the a-priori knowledge of the average number of positive labels per sample (see Tables I and II).
-  • General label noise methods: From this category, we considered the ELR approach adapted as in [10], the LL family of approaches (i.e., LL-R, LL-Ct, and $L$-Cp) [16], and the BoostLU [19] in combination with $L$-Ct.
-  • PUL methods: We considered a single SOTA method from this category, i.e., Dist-PU [39], adapted for SPML by simply applying it to each class independently.
-  • SPML-specific methods: The main comparison methods we considered for SPML are ROLE [12], GR Loss [18], LAGC [17], and Mutual label enhancement for sIngle-positive Multi-label lEarning (MIME) [20].
-  • Upper bound methods: For analytical comparison with GT training, we considered two cases: i) training with GT multi-labels using a BCE loss (GT w/ BCE), and ii) the with true Negatives (IUN) approach [12], where in addition to the single positive labels, all true negatives are provided for training.
-IUN allows us to assess the performance in a hypothetical scenario where all true negatives have been retrieved.
-False negatives are ignored during training, meaning that the model is still provided with a single positive label for each image.
-IV-C Results
 TABLE III: Test Performance Comparison of Different Methods on the reBEN-Random Dataset.
@@
 The Related Standard Deviations Are Reported in Brackets
-mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
+$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
 Ground-truth Training GT w/ BCE $71.75(0.04)$ 4.264 (0.003) 3.428 (0.005) 93.09 (0.01) 66.63 (0.12) 72.07 (0.02) 63.24 (0.35)
@@
 MIME $58.64(0.35)$ 4.894 (0.017) 5.632 (0.064) 90.66 (0.13) 55.13 (0.36) 62.40 (0.46) 50.39 (0.21)
-Proposed AdaGC 68.18 (0.35) 4.513 (0.012) 4.073 (0.033) 92.52 (0.02) 64.65 (0.37) 70.36 (0.84) 63.37 (0.82)
+RS-specific Methods RCML $64.91(0.12)$ 4.585 (0.005) 4.467 (0.020) 91.24 (0.04) 52.90 (0.32) 78.57 (1.29) 42.42 (0.34)
+LCR $54.56(0.32)$ 4.864 (0.009) 5.912 (0.044) 87.16 (0.03) 28.52 (0.14) 73.82 (1.81) 19.31 (0.08)
+Proposed AdaGC $67.85(0.05)$ 4.553 (0.012) 4.165 (0.033) 92.49 (0.01) 64.46 (0.24) 70.43 (0.59) 63.22 (0.10)
 a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
+IV-B Experimental Setup
+We adopt ResNet-34 and ResNet-50 as backbones for the reBEN and AID datasets, with batch sizes of 512 and 32, respectively.
+ResNet-34 is chosen for reBEN as an efficient compromise, since our focus is on evaluating method effectiveness rather than maximizing absolute accuracy.
+For AID, we use ImageNet-pretrained weights given its RGB input.
+To adhere to a realistic weakly supervised scenario where ground-truth multi-labels are unavailable, we strictly avoid the use of clean data for model selection.
+Consequently, all hyperparameters are tuned via grid search on the noisy validation set, simulating the absence of clean labels.
+While this strategy relies on imperfect supervision, it prevents the introduction of “oracle” information and ensures that the reported performance reflects the method’s capability to generalize without privileged access to clean labels.
+Despite the use of noisy validation labels, the resulting metrics remain informative for model selection.
+Under SPML noise, random errors are inherently difficult for deep models to memorize, and thus the validation signal largely reflects the underlying clean structure.
+As confirmed in our experiments, noisy mAP validation scores maintain a strong correlation with clean test performance, making them a practical and reliable indicator in the absence of ground-truth labels.
+To reduce computational cost, tuning is performed on subsets (20% for reBEN, 50% for AID) and fewer epochs (10) for comparison methods.
+For AdaGC, the Mixup Beta parameter $\alpha$ is set to 1.
+The GC loss weight $\lambda$ is set to 3, 5, 9, and 8 for reBEN-Random, reBEN-Dominant, AID-Random and AID-Manual, respectively, and the learning rate is set to 1e-3 and 5e-4 for reBEN and AID, respectively, based on tuning results.
+For the final evaluation, all models are trained for 70 epochs on the full training sets using the optimal hyperparameters and evaluated on the clean test sets.
+To ensure robustness and statistical reliability, each experiment is repeated for three Monte Carlo runs.
+We report the mean and unbiased standard deviation of all metrics.
 TABLE IV: Test Performance Comparison of Different Methods on the reBEN-Dominant Dataset.
@@
 The Related Standard Deviations Are Reported in Brackets
-mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
+$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
 Ground-truth Training GT w/ BCE $71.75(0.04)$ 4.264 (0.003) 3.428 (0.005) 93.09 (0.01) 66.63 (0.12) 72.07 (0.02) 63.24 (0.35)
@@
 MIME $55.61(0.07)$ 6.417 (0.061) 8.982 (0.125) 89.90 (0.03) 43.54 (0.39) 76.38 (1.19) 33.36 (0.55)
-Proposed AdaGC 60.22 (0.24) 5.618 (0.004) 6.926 (0.016) 90.15 (0.05) 58.78 (0.44) 61.28 (0.98) 62.48 (0.62)
+RS-specific Methods RCML $58.26(0.82)$ 5.890 (0.134) 8.035 (1.181) 88.73 (1.05) 40.75 (1.27) 83.06 (2.35) 30.93 (2.88)
+LCR $55.64(0.20)$ 6.579 (0.045) 8.151 (0.080) 88.67 (0.00) 34.25 (0.06) 83.24 (0.54) 24.05 (0.04)
+Proposed AdaGC $60.06(0.16)$ 5.581 (0.026) 6.877 (0.054) 90.00 (0.06) 59.08 (0.25) 60.19 (0.47) 63.43 (0.37)
 a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
-This section presents a detailed analysis of the experimental results reported in Tables III, IV, and V, comparing the performance of AdaGC against various baselines, PUL, SPML-specific, and general label noise methods across the reBEN and AID datasets under different SPML noise types.
-Several general trends are observable across all experimental scenarios.
-Our proposed AdaGC consistently achieves the best mAP values across all three evaluated scenarios.
-Furthermore, it demonstrates SOTA SPML performance in terms of coverage, rankloss, and mF1, highlighting its overall effectiveness and robustness in handling SPML learning for RS imagery.
-The upper-bound method IUN, which leverages knowledge of true negatives while only observing single positives, consistently achieves a good mAP.
-However, due to its bias towards single positive labels, it typically exhibits high mprecision but comparatively lower mrecall, which consequently impacts its mF1.
-The AN and AN-LS baselines are particularly sensitive to the inherent bias introduced by single positive labels and the prevalence of false negatives.
-This sensitivity is reflected in their high mprecision but significantly low mrecall, leading to a generally low mAP across all datasets.
-In contrast, the WAN method demonstrates greater robustness compared to AN and AN-LS.
-It achieves a better balance between mprecision and mrecall, making it the best-performing baseline method in terms of mAP across the considered scenarios.
-EPR employs a distinct strategy by training solely on single positive samples and incorporating a regularization term to enforce the prediction of an average number of positive labels.
-This approach consistently results in EPR being the model with the highest mrecall across all datasets and noise types.
-However, given its reliance only on positive supervision, its mprecision is generally lower compared to methods that attempt to identify true negatives.
-Additionally, the consistently high mrecall scores of EPR are attributable to its unique reliance on a-priori knowledge of the average number of positive labels per image, a piece of information not utilized by other methods under comparison.
-Next subsections discuss the results obtained on each specific dataset and scenario.
+In addition to mAP, we report two ranking-based metrics, i.e., coverage and ranking loss (Rankloss), and four accuracy-based metrics, i.e., overall accuracy (OA), mean F1 score (mF1), mean precision (mprecision), and mean recall (mrecall).
+The metrics are defined as follows.
+  • Coverage measures how far we need to go down the ranked label list to cover all true labels:
+    $\frac{1}{N}\sum_{i=1}^{N}\max_{y\in\mathbf{y}_{i}}\text{rank}_{f}(\mathbf{x}_{i},y)-1,$ (22)
+    where $\text{rank}_{f}(x_{i},y)$ is the position of the ground-truth (GT) label $y_{i,c}$ in the ranking list predicted by $f$;
+  • Ranking loss computes the average fraction of label pairs that are incorrectly ordered:
+    $\small\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathbf{y}_{i}||\overline{\mathbf{y}_{i}}|}\sum_{(y_{1},y_{0})\in\mathbf{y}_{i}\times\overline{\mathbf{y}_{i}}}\mathbb{I}(f(\mathbf{x}_{i},y_{1})\leq f(\mathbf{x}_{i},y_{0})),$ (23)
+    where $\mathbf{y}_{i}$ and $\overline{\mathbf{y}_{i}}$ are the sets of relevant (positive) and irrelevant (negative) labels, respectively, and $\mathbb{I}(\texttt{True})=1$ otherwise 0;
+  • OA measures the percentage of samples with all labels correctly predicted:
+    $\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}\mathbb{I}(\hat{\mathbf{y}}_{i,c}=\mathbf{y}_{i,c});$ (24)
+  • mF1 is the average of F1-scores computed per class:
+    $\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Prec}_{c}\cdot\text{Rec}_{c}}{\text{Prec}_{c}+\text{Rec}_{c}}.$ (25)
+We compared AdaGC against a wide array of existing methods, categorized as follows.
+  • Baseline methods: We considered several baselines [Cole_2021_CVPR], such as AN, AN-LS, Weak (WAN), and EPR.
+All use a BCE loss at their core, while EPR uses only the BCE positive term and adds a regularization term, assuming the a-priori knowledge of the average number of positive labels per sample (see Tables I and II).
+  • General label noise methods: From this category, we considered the ELR approach adapted as in [Burgert_2022_TGRS], the LL family of approaches (i.e., LL-R, LL-Ct, and LL-Cp) [Kim_2022_CVPR], and the BoostLU [Kim_2023_CVPR] in combination with LL-Ct.
+  • PUL methods: We considered a single state-of-the-art (SOTA) method from this category, i.e., Dist-PU [Zhao_2022_CVPR], adapted for SPML by simply applying it to each class independently.
+  • SPML-specific methods: The main comparison methods we considered for SPML are ROLE [Cole_2021_CVPR], GR Loss [Chen_2024_IJCAI], LAGC [Xie_2022_NIPS], and Mutual label enhancement for sIngle-positive Multi-label lEarning (MIME) [Liu_2023_ICML].
+  • Upper bound methods: For analytical comparison with GT training, we considered two cases: i) training with GT multi-labels using a BCE loss (GT w/ BCE), and ii) the with true Negatives (IUN) approach [Cole_2021_CVPR], where in addition to the single positive labels, all true negatives are provided for training.
+IUN allows us to assess the performance in a hypothetical scenario where all true negatives have been retrieved.
+False negatives are ignored during training, meaning that the model is still provided with a single positive label for each image.
+  • RS-specific methods: We also included RCML [Aksoy_2024_TNNLS] and LCR [hua_learning_2020], which were originally developed for general noisy multi-label RS classification.
+RCML leverages a collaborative learning framework with a ranking loss for sample selection to mitigate label noise, while LCR uses word embedding similarity to regularize model training under noisy annotations.
+To adapt them to SPML, we use only the missing-class labels term in RCML and perform selection per class rather than per sample.
+For LCR, we use RemoteCLIP’s [liu_remoteclip_2024] text encoder to encode multi-word class names.
+These adaptations ensure a fair comparison while preserving the core noise-robust strategies of the original methods.
+IV-C Benchmark Results
+This section analyzes the benchmark results reported in Tables IIIVI.
+Across all scenarios, AdaGC consistently achieves the highest mAP and strong SPML performance in terms of coverage, rankloss, and mF1, demonstrating its effectiveness and robustness in handling SPML in RS imagery.
+The upper-bound IUN performs well in mAP but suffers from low mrecall due to its inherent single-positive bias.
+AN and AN-LS are highly sensitive to false negatives, exhibiting high mprecision but low mrecall and consequently poor mAP.
+In contrast, WAN is more robust and achieves the best baseline mAP by better balancing precision-recall trade-offs.
+EPR, trained exclusively on positive labels with a cardinality regularizer, consistently shows the highest mrecall across datasets but low mprecision, reflecting its tendency to overestimate positives.
+Its advantage largely stems from the use of the true average number of labels per image, a prior not shared by other competitors.
+The two RS-specific methods designed for general label noise exhibit clear limitations under SPML conditions.
+The following subsections detail dataset- and scenario-specific observations.
 IV-C1 reBEN-Random
-TABLE V: Test Performance Comparison of Different Methods on the AID Dataset.
+TABLE V: Test Performance Comparison of Different Methods on the AID-Random Dataset.
 The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic.
 The Related Standard Deviations Are Reported in Brackets
-mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
-Ground-truth Training GT w/ BCE 88.90 (0.09) 5.784 (0.036) 1.649 (0.177) 94.35 (0.14) 83.70 (0.34) 87.43 (1.02) 81.91 (0.63)
-IUN 82.85 (0.33) 6.311 (0.060) 2.741 (0.087) 91.21 (0.37) 71.17 (2.28) 95.04 (1.41) 60.64 (2.15)
-Baseline Methods AN 65.41 (1.28) 8.403 (0.177) 10.908 (0.349) 71.74 (0.19) 27.16 (0.12) 86.42 (2.26) 16.92 (0.09)
-AN-LS 64.89 (0.71) 11.175 (0.189) 21.113 (1.477) 70.81 (0.25) 20.11 (0.57) 80.53 (5.82) 12.22 (0.44)
-WAN 68.70 (0.49) 7.481 (0.185) 7.575 (0.647) 78.36 (0.17) 50.14 (0.47) 84.57 (1.09) 36.95 (0.35)
-EPRa 68.62 (1.87) 7.277 (0.093) 8.453 (0.390) 84.91 (0.57) 70.09 (1.12) 62.33 (0.85) 81.76 (1.78)
-General Label Noise Methods ELR 69.97 (0.69) 7.176 (0.014) 5.401 (0.052) 88.25 (0.20) 59.36 (0.89) 74.95 (0.96) 56.19 (0.35)
-$L$-Cp 64.71 (1.07) 7.629 (0.132) 7.426 (0.465) 78.73 (0.57) 45.64 (1.16) 76.96 (3.89) 36.35 (0.72)
-$L$-Ct 67.48 (1.53) 8.302 (0.260) 6.846 (0.374) 85.78 (0.65) 55.78 (0.91) 78.65 (4.73) 48.19 (0.88)
-$L$-R 69.56 (0.43) 8.314 (0.023) 7.958 (0.313) 80.75 (0.58) 51.42 (2.25) 85.28 (2.87) 39.15 (2.38)
-BoostLU 68.99 (1.11) 7.218 (0.113) 5.733 (0.308) 85.79 (0.94) 54.86 (0.83) 79.39 (1.79) 47.40 (0.97)
-PUL Methods Dist-PUa 71.38 (0.28) 6.493 (0.112) 3.761 (0.177) 90.02 (0.11) 60.59 (3.33) 78.60 (6.94) 57.09 (3.12)
-SPML Methods ROLE 56.68 (0.79) 13.238 (0.053) 19.038 (0.294) 76.44 (0.38) 43.66 (0.64) 84.18 (3.33) 31.10 (0.49)
-GR 62.69 (0.79) 11.053 (0.091) 20.911 (0.577) 72.34 (0.08) 28.62 (0.57) 83.20 (4.34) 18.36 (0.31)
-LAC 62.61 (1.01) 7.746 (0.086) 9.187 (0.231) 72.74 (0.03) 28.84 (1.87) 82.41 (4.61) 18.56 (1.31)
-MIME 61.96 (1.94) 7.150 (0.166) 7.016 (0.799) 73.94 (0.28) 31.96 (1.03) 81.02 (2.73) 20.82 (0.66)
-Proposed AdaGC 74.82 (0.90) 6.607 (0.054) 4.381 (0.154) 86.51 (0.49) 60.68 (2.24) 77.86 (1.33) 60.03 (2.15)
+$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
+Ground-truth Training GT w/ BCE $89.01(0.32)$ 5.834 (0.011) 1.657 (0.061) 94.29 (0.19) 83.12 (0.66) 88.87 (0.38) 79.58 (1.28)
+IUN $82.24(0.61)$ 6.622 (0.031) 3.408 (0.026) 90.88 (0.19) 72.65 (0.33) 93.13 (0.86) 62.87 (0.49)
+Baseline Methods AN $64.89(1.28)$ 8.557 (0.164) 11.360 (0.351) 71.60 (0.23) 26.83 (0.18) 86.05 (2.18) 16.71 (0.13)
+AN-LS $64.40(0.73)$ 11.347 (0.231) 21.664 (1.657) 70.65 (0.28) 19.97 (0.63) 79.98 (5.71) 12.14 (0.47)
+WAN $68.12(0.46)$ 7.620 (0.173) 8.055 (0.588) 78.06 (0.16) 49.53 (0.42) 84.19 (1.05) 36.45 (0.27)
+EPRa $68.38(1.84)$ 7.395 (0.079) 9.086 (0.412) 84.76 (0.57) 69.82 (1.10) 62.30 (0.85) 81.11 (1.68)
+General Label Noise Methods ELR $69.13(0.73)$ 7.469 (0.063) 5.914 (0.152) 87.99 (0.12) 59.04 (0.84) 74.80 (1.00) 55.91 (0.30)
+$L$-Cp $64.42(1.06)$ 7.792 (0.141) 7.931 (0.471) 78.64 (0.58) 45.47 (1.25) 76.83 (3.94) 36.25 (0.78)
+$L$-Ct $66.96(1.60)$ 8.574 (0.243) 7.647 (0.322) 85.42 (0.65) 55.41 (0.91) 78.43 (4.80) 47.92 (0.75)
+$L$-R $68.85(0.44)$ 8.526 (0.018) 8.616 (0.294) 80.39 (0.58) 50.89 (2.19) 84.92 (2.85) 38.75 (2.31)
+BoostLU $68.00(0.97)$ 8.104 (0.127) 6.656 (0.519) 85.44 (0.97) 54.48 (0.86) 79.19 (1.90) 47.14 (1.00)
+PUL Methods Dist-PUa $70.85(0.47)$ 6.692 (0.133) 4.355 (0.265) 89.86 (0.09) 60.48 (3.26) 78.55 (6.93) 57.04 (3.10)
+SPML Methods ROLE $56.35(0.77)$ 13.581 (0.041) 19.800 (0.282) 76.10 (0.38) 43.26 (0.73) 83.73 (3.32) 30.84 (0.57)
+GR $62.14(0.73)$ 11.236 (0.090) 21.537 (0.567) 72.11 (0.09) 28.32 (0.56) 82.59 (4.41) 18.18 (0.32)
+LAC $62.27(0.92)$ 7.902 (0.091) 9.674 (0.266) 72.57 (0.01) 28.59 (1.81) 82.17 (4.10) 18.38 (1.22)
+MIME $61.55(1.86)$ 7.308 (0.175) 7.474 (0.756) 73.67 (0.26) 31.55 (1.32) 81.05 (3.42) 20.52 (0.86)
+RS-specific Methods RCML $64.24(1.41)$ 7.522 (0.107) 8.338 (0.445) 80.10 (0.64) 45.45 (0.47) 75.62 (4.66) 37.84 (0.23)
+LCR $62.62(0.47)$ 8.269 (0.165) 10.453 (0.500) 71.68 (0.13) 23.48 (1.01) 78.12 (2.50) 15.06 (0.58)
+Proposed AdaGC $75.67(0.26)$ 6.862 (0.209) 5.636 (1.303) 83.39 (2.40) 64.16 (1.57) 68.89 (5.58) 71.82 (6.80)
 a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
-Table III summarizes the performance of all methods on the reBEN dataset when subjected to Random SPML noise.
-Our proposed AdaGC demonstrates superior overall performance, achieving the best mAP, OA, mF1, coverage, and rankloss scores.
-While it does not achieve the best performance in mprecision and mrecall, AdaGC shows the best balance between two, as reflected in the mF1 score.
-General label noise and PUL methods show good performance in this scenario, outperforming SPML-specific methods.
-Additionally, WAN demonstrates to be a strong baseline, outperforming all other baselines as well as many other comparison methods.
-The reason might be found in the hyperparameter selection.
-Indeed, we tuned them for each method individually by monitoring the noisy validation mAP.
-While noisy mAP can be a good indicator for early learning detection when evaluated continuously during training, its fluctuations in hyperparameters selection, combined with methodologies that stem from AN-like approaches (e.g., pseudo-labelling or label correction approaches), can result in suboptimal hyperparameters’ selection.
-However, we argue that methods benchmarking should reflect real-world settings as much as possible, which is ensured in our experimental setting.
-Comparing the results with GT training, we can observe that the AdaGC approaches the performance of training with GT multi-labels, especially in terms of mF1.
-Compared to IUN, AdaGC achieves a worse mAP score and a better mF1 score.
-Although false negatives do not affect IUN in mAP results, the mF1 scores emphasize the importance of methodologies that leverage pseudo-labels or label correction.
-Such approaches are vital for accurately reconstructing multi-labels, thus preventing classifier confidence bias and achieving higher mF1 scores.
+Table III presents the results on the reBEN dataset under Random SPML noise.
+AdaGC delivers the strongest overall performance, achieving the highest mAP, OA, and mF1, together with the best coverage and rankloss scores.
+Although not optimal in mprecision or mrecall individually, it provides the most effective precision–recall trade-off, as reflected by its mF1.
+General label-noise and PUL methods perform well in this setting, often surpassing SPML-specific approaches.
+WAN remains particularly competitive among baselines.
+Within RS-specific methods, RCML offers noticeable improvements compared to baselines, whereas LCR contributes little due to its fixed class-correlation assumptions.
+Compared with GT training, AdaGC closely approaches the full-label upper bound, especially in mF1.
+Compared to IUN, AdaGC attains slightly lower mAP but higher mF1.
+This highlights the importance of incorporating pseudo-labeling or label-correction mechanisms to recover missing labels and avoid confirmation bias.
 IV-C2 reBEN-Dominant
-While Random SPML can be a good simulation for cases where both real-world single labels and GT multi-labels are difficult to gather, it provides a simplified scenario.
-Indeed, random single labels can still provide all the necessary information for describing the land cover, since the single labels are chosen randomly among the GT labels.
-On the contrary, dominant single labels are completely driven by the actual observed image, leading to cases where specific non-dominant class patterns are entirely missing from the training set, making this scenario more challenging.
-This can be observed in Table IV, where all methods show lower absolute performance compared to their counterparts in Table III.
-Despite this increased difficulty, the proposed AdaGC maintains its leading position, achieving the highest mAP, mF1, coverage, and rankloss scores.
-Instead, the comparison methods struggle to outperform the WAN and EPR baseline methods.
-WAN shows the second best mAP score, and its mF1 score surpasses all SPML-specific methods.
-Only methods based on the LL family and Dist-PU are able to outperform WAN in terms of mF1, struggling against EPR.
-Due to the regularization term that affects the number of predicted positive labels, EPR can overcome the strong bias introduced by Dominant SPML in terms of mF1 score, despite its low mAP score.
-IV-C3 AID
+Random SPML provides a simplified single-positive noise setting, whereas Dominant SPML is a more practical scenario driven by the visually dominant class in each image, resulting in systematic omission of non-dominant classes.
+This significantly increases task difficulty, as reflected in Table IV, where all methods exhibit lower absolute performance compared with Table III.
+Despite this challenge, AdaGC remains the strongest method, achieving the best mAP, mF1, coverage, and rankloss.
+In contrast, most competing approaches fail to surpass the WAN and EPR baselines.
+WAN obtains the second-best mAP and outperforms all SPML-specific methods in mF1.
+LL-based methods and Dist-PU offer additional gains over WAN in mF1, yet still fall short of EPR.
+Thanks to its cardinality regularization, EPR effectively mitigates the strong bias induced by Dominant SPML, yielding competitive mF1 despite its low mAP.
+RCML remains relatively strong but continues to underperform the proposed AdaGC.
+IV-C3 AID-Random
 (a) reBEN
@@
 Figure 4: Histograms of the number of GT labels per image in the training sets.
-Table V displays the performance of all methods on the AID dataset with Random SPML noise.
-This dataset, benefiting from ImageNet pre-training and richer spatial information, generally shows higher absolute performance metrics across the board compared to reBEN, despite its generally higher flip rates $\beta_{c}$.
-Consistent with its overall very good performance, AdaGC emerges at the top of the ranking for most metrics.
-It achieves the best mAP score, and the second best mF1, mrecall, coverage, and rankloss scores.
-Notably, Dist-PU shows outstanding performance, on par with AdaGC’s mF1, and achieves the best OA, coverage, and rankloss scores.
-Additionally, EPR shows the best mF1 score.
-Nevertheless, the top performance of Dist-PU and EPR can be explained in terms of prior information available.
-Indeed, both methods exploit some a-priori information regarding the GT multi-labels, which is not the case for the proposed AdaGC.
-Dist-PU assumes that the knowledge of the prior distributions of all classes is available, and exploits this information to match it with the class priors of the predicted labels, while EPR uses the knowledge of the expected number of positive labels per image.
-This information is particularly effective on the AID dataset, as it shows, in general, a higher number of labels for each image, as shown in Fig.
-4.
-This makes EPR exhibit the best mrecall among all methods, on par with GT training, which is high enough to counterbalance the worst mprecision across the board, which highlights a label overestimation tendency.
+Table V reports the results on the AID dataset under Random SPML noise.
+Owing to ImageNet pre-training and richer spatial information, this dataset yields higher absolute performance than reBEN, despite larger flip rates $\beta_{c}$.
+AdaGC shows the best performance across all metrics.
+It achieves the highest mAP and the second-best mF1, mrecall, coverage, and rankloss scores.
+Dist-PU also shows excellent results, matching AdaGC in mF1 while achieving the best OA, coverage, and rankloss.
+Meanwhile, EPR reaches the highest mF1 overall.
+The strong outcomes of Dist-PU and EPR are largely attributable to their use of prior information about the underlying GT label distributions, which is not available to AdaGC.
+Dist-PU assumes knowledge of class-prior distributions and aligns predicted label frequencies accordingly, whereas EPR relies on the expected number of positive labels per image.
+These priors are particularly beneficial for AID, which contains more labels per image on average (Fig.
+4).
+As a result, EPR achieves the highest mrecall, although this comes at the cost of the lowest mprecision, reflecting its tendency to overestimate positives.
+Still, RCML achieves some comparable results to other considered methods, whereas the improvement brought by LCR remains limited.
+IV-C4 AID-Manual
+TABLE VI: Test Performance Comparison of Different Methods on the AID-Manual Dataset.
+The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic.
+The Related Standard Deviations Are Reported in Brackets
+$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
+Ground-truth Training GT w/ BCE $89.01(0.32)$ 5.834 (0.011) 1.657 (0.061) 94.29 (0.19) 83.12 (0.66) 88.87 (0.38) 79.58 (1.28)
+IUN $82.71(0.60)$ 6.638 (0.085) 3.068 (0.191) 91.47 (0.16) 74.83 (0.42) 94.18 (3.95) 65.53 (0.89)
+Baseline Methods AN $67.66(0.73)$ 8.589 (0.140) 11.276 (0.283) 71.30 (0.08) 29.01 (0.59) 83.94 (1.69) 18.83 (0.56)
+AN-LS $66.75(0.62)$ 11.556 (0.061) 23.300 (0.640) 71.34 (0.17) 28.41 (1.81) 84.61 (1.94) 18.32 (1.19)
+WAN $70.28(0.50)$ 7.721 (0.103) 8.364 (0.276) 75.35 (0.29) 45.28 (1.77) 88.30 (4.71) 32.21 (1.68)
+EPRa $71.92(1.36)$ 7.251 (0.118) 8.766 (0.358) 84.89 (0.50) 71.30 (1.06) 63.86 (1.49) 83.84 (0.56)
+General Label Noise Methods ELR $76.99(0.51)$ 7.278 (0.107) 5.353 (0.232) 88.26 (0.31) 69.09 (0.63) 77.43 (0.47) 68.55 (1.33)
+$L$-Cp $69.65(0.54)$ 7.854 (0.073) 8.982 (0.470) 70.87 (0.07) 23.14 (0.23) 80.08 (3.81) 14.94 (0.28)
+$L$-Ct $72.46(0.11)$ 7.564 (0.104) 7.788 (0.525) 76.04 (0.12) 42.70 (0.69) 85.52 (2.13) 30.93 (0.24)
+$L$-R $73.89(0.29)$ 7.743 (0.007) 6.560 (0.094) 83.95 (0.05) 63.51 (1.55) 80.53 (1.29) 54.60 (1.90)
+BoostLU $71.35(1.77)$ 8.441 (0.061) 8.976 (0.247) 75.61 (0.06) 42.58 (0.40) 83.31 (6.37) 31.62 (0.56)
+PUL Methods Dist-PUa $74.90(1.88)$ 6.574 (0.021) 3.652 (0.085) 90.23 (0.34) 66.77 (1.16) 80.98 (0.23) 62.70 (1.32)
+SPML Methods ROLE $64.36(1.99)$ 13.253 (0.101) 17.429 (0.973) 76.41 (0.64) 49.11 (2.03) 86.08 (4.40) 36.30 (2.01)
+GR $67.18(0.65)$ 11.042 (0.191) 21.481 (0.353) 71.63 (0.14) 29.88 (1.34) 86.64 (4.63) 19.93 (1.17)
+LAC $65.80(0.81)$ 7.878 (0.066) 9.395 (0.259) 72.64 (0.16) 31.72 (1.58) 80.81 (4.74) 21.03 (1.04)
+MIME $65.39(1.07)$ 7.240 (0.147) 6.333 (0.424) 73.10 (0.14) 32.62 (0.89) 84.73 (6.70) 21.70 (0.65)
+RS-specific Methods RCML $72.87(0.57)$ 7.453 (0.019) 7.393 (0.266) 80.46 (1.44) 52.99 (3.67) 85.83 (2.34) 43.08 (5.79)
+LCR $66.61(0,96)$ 8.322 (0.078) 10.669 (0.472) 71.79 (0.22) 31.83 (1.46) 85.52 (1.57) 21.03 (1.02)
+Proposed AdaGC $79.07(0.91)$ 6.673 (0.103) 4.620 (0.558) 82.66 (1.01) 71.95 (0.94) 68.25 (2.35) 82.61 (1.28)
+a EPR and Dist-PU use additional a-priori information, which is assumed unknown in all other approaches.
+Table VI reports the results on the AID dataset using Manual SPML labels to further evaluate AdaGC’s practical effectiveness.
+Benefiting from human-induced bias in single-positive labels, most methods achieve higher performance compared with AID-Random.
+In this setting, AdaGC remains highly effective, getting the highest mAP and mF1, and ranking among the top methods across nearly all metrics.
+Dist-PU again achieves the best OA, coverage, and rankloss by leveraging extra prior information.
+RS-specific methods show similar trends as before.
+RCML provides moderate improvements over other SPML or label-noise baselines, while LCR contributes little due to the fixed class-correlation assumptions.
+Overall, the AID-Manual results confirm that higher-quality single-positive labels enhance learning, and AdaGC consistently delivers robust performance without relying on additional priors.
 IV-D Ablation Experiments
@@
 Without making specific claims, we report the accuracies of the teacher models below.
-IV-D1 Early Learning Detection
-(a) Random-mAP
-(b) Random-mF1
-(c) Dominant-mAP
-(d) Dominant-mF1
-Figure 5: Validation accuracies (mAP and mF1) versus training time (epoch) obtained with predefined warm-up length (10, 20, 30) and our proposed early learning detection strategy (AdaGC).
+IV-D1 Impact of Early Learning Detection
+(a) Random: mAP
+(b) Random: mF1
+(c) Dominant: mAP
+(d) Dominant: mF1
+Figure 5: Validation accuracies (mAP and mF1) over training epochs obtained with predefined warm-up lengths (10, 20, 30) and our proposed early learning detection strategy (AdaGC).
 We evaluate the effectiveness of early learning detection for adaptively triggering GC by comparing AdaGC with variants that use manually predefined trigger points.
 Fig.
-5 plots their true validation accuracies (with respect to the GT labels) over training time.
-As shown, the proposed adaptive detection strategy successfully activates GC at an appropriate point, resulting in the most significant performance improvements.
+5 plots their validation accuracies (wrt GT labels) over training time.
+The proposed adaptive detection strategy successfully activates GC at an appropriate point, resulting in the most significant performance improvements.
+This demonstrates that the timing of GC activation is essential for model convergence stability.
 In contrast, triggering GC too early or too late leads to suboptimal results.
 Notably, activating GC too late leads to even worse performance, highlighting the detrimental effect of overfitting to label noise.
-Furthermore, Table VI reports the final test accuracies.
-AdaGC achieves competitive or comparable mAP and mF1 scores compared to those obtained by using fixed warm-up lengths, indicating the robustness of the proposed early learning detection strategy.
-TABLE VI: Accuracies (%) on the Test Set Obtained With Different GC Trigger Settings After 70 Epochs
+Furthermore, we report their final test accuracies in Table VII.
+AdaGC achieves competitive or comparable mAP and mF1 scores compared to those obtained using fixed warm-up lengths, indicating the robustness of the proposed early learning detection strategy.
+TABLE VII: Accuracies (%) on the Test Set Obtained With Different GC Trigger Settings After 70 Epochs
 Random Dominant
@@
 Trigger after 30 epochs 63.18 54.77 58.62 55.71
-IV-D2 Teacher and Student Models
-We examine the roles of the student and teacher models in two modules, i.e., early learning detection and pseudo-label generation.
-We first illustrate the smoothness of the noisy validation accuracies produced by the student and teacher models in Fig.
+We demonstrate the necessity of using the teacher model for early learning detection by comparing the smoothness of noisy validation accuracies from the student and teacher models in Fig.
 6.
-The accuracy curves from both models exhibit similar trends.
-Nevertheless, those from the teacher models are noticeably smoother than those from the student models.
-This makes teacher models more suitable for early learning detection.
-In contrast, using student models for detection is unstable, as the detection is more likely to fall into a local maximum, leading to suboptimal results.
+Although both models exhibit similar overall trends, the teacher model produces noticeably smoother accuracy curves.
+This stability makes it more suitable for reliable early learning detection.
+In contrast, detections based on the student model are more prone to local maxima, resulting in unstable and suboptimal outcomes.
 (a) Random
@@
 Figure 6: Noisy validation mAP (with respect to noisy labels) versus training time (epoch) obtained by AN on the reBEN dataset.
-TABLE VII: Accuracies (%) on the Test Set Obtained With Pseudo Labels
-Generated by Different Models, Where T and S Represent
-Teacher and Student Models, Respectively
-Random Dominant
+IV-D2 Effect of the Dual-EMA Mechanism
+We investigate the effect of the proposed dual-EMA mechanism by comparing it with single-EMA and no-EMA settings, as shown in Table VIII.
+The dual-EMA strategy consistently achieves the most robust performance by striking a balance between stability and adaptability.
+In contrast, the single-EMA and no-EMA variants exhibit suboptimal results, particularly when the EMA smoothing on student predictions is removed.
+In both Random and Dominant noise settings, the EMA-smoothed student predictions are more stable than the teacher-only ones, owing to the teacher’s slower update rate.
+Nevertheless, without EMA smoothing (no-EMA), the student’s raw predictions tend to amplify errors, resulting in additional performance degradation.
+This trend is further confirmed in Fig.
+7, where pseudo-label quality is shown to have a strong impact on AdaGC’s performance.
+As the pseudo-label quality deteriorates, performance drops markedly, especially for the teacher-only and no-EMA cases.
+The dual-EMA mechanism mitigates these issues by smoothing student predictions and integrating teacher outputs, thereby generating more reliable pseudo labels.
+Considering the varying update dynamics of the student and teacher models during training, a more adaptive combination strategy beyond the current half–half scheme is expected to enhance robustness, which we leave for future exploration.
+TABLE VIII: Test set accuracies (%) obtained using different pseudo-label generation strategies, where T, S, and S denote teacher model predictions, EMA-smoothed student model predictions, and student model predictions without EMA, respectively
+EMA Settings Random Dominant
 mAP $\uparrow$ mF1 $\uparrow$ mAP $\uparrow$ mF1 $\uparrow$
-AdaGC (T+S, $\gamma=0.5$) 68.18 64.65 60.22 58.78
-Only T ($\gamma=1$) 61.74 63.49 42.01 46.40
-Only S ($\gamma=0$) 67.93 61.52 61.11 57.42
-T+S (S w/o EMA, $\gamma=0.5$) 53.78 61.94 27.50 35.71
-We then investigate the roles of the student and teacher models in pseudo-label generation.
-As shown in Table VII, combining predictions from both models yields the most robust performance.
-In the Random noise setting, relying solely on either the student or the teacher model fails to achieve optimal mAP and mF1 scores.
-In the Dominant noise setting, using only the student model produces more stable results than using only the teacher model.
-This is partly due to the teacher model’s relatively slow update rate, which lags behind the student model’s more responsive prediction updates.
-As a result, the teacher may fail to adapt quickly enough to changing training dynamics, making it more vulnerable to overfitting to the dominant noisy labels.
-In such scenarios, swift and timely pseudo-label updates are essential for maintaining performance and robustness.
-Nevertheless, combining predictions from both models offers a balance between stability and adaptability, ensuring robust performance across different types of label noise.
-Finally, we evaluate the role of applying EMA to student predictions by replacing them with the student model’s direct online predictions.
-The results, reported in the last row of Table VII, show a significant performance drop in both cases, with a more sever decrease in the Dominant setting.
-This highlights the importance of using EMA to stabilize the student predictions and ensure higher-quality pseudo-labels in GC.
+dual AdaGC 68.18 64.65 60.22 58.78
+single Only T ($\gamma=1$) 61.74 63.49 42.01 46.40
+Only S ($\gamma=0$) 67.93 61.52 60.11 57.42
+T+S ($\gamma=0.5$) 53.78 61.94 27.50 35.71
+no Only S ($\gamma=0$) 49.30 52.89 16.41 25.24
+(a) Random: pseudo-label mAP
+(b) Random: pseudo-label mF1
+(c) Random: validation mAP
+(d) Random: validation mF1
+(e) Dominant: pseudo-label mAP
+(f) Dominant: pseudo-label mF1
+(g) Dominant: validation mAP
+(h) Dominant: validation mF1
+Figure 7: Pseudo-label quality (a), (b), (e), (f) and AdaGC validation accuracies (mAP and mF1) (c), (d), (g), (h) over training epochs using different pseudo-label generation strategies, where T, S, and S denote teacher model predictions, EMA-smoothed student model predictions, and student model predictions without EMA, respectively.
+(a)-(d) Random single-positive labels case.
+(e)-(h) Dominant single-positive labels case.
 IV-D3 Role of Mixup
-We remove Mixup from AdaGC and report the results in Table VIII to assess its impact.
-As shown, the absence of Mixup leads to a consistent drop in performance across both noise settings, particularly in the Random case, where mAP and mF1 decrease by approximately 2%.
-These findings suggest that Mixup enhances generalization, especially under random noise, by encouraging smoother decision boundaries.
-Although its effect is less pronounced under Dominant noise, it still provides measurable benefits, underscoring its complementary role in improving the robustness of AdaGC.
-TABLE VIII: Accuracies (%) on the Test Set Obtained With and Without Mixup
+We then investigate the effect of Mixup under both Random and Dominant noise settings.
+As shown in Table IX, removing Mixup consistently reduces performance, particularly in the Random case, with about a 2% drop in mAP and mF1.
+This indicates that Mixup enhances generalization by enforcing smoother decision boundaries.
+It can complement the GC mechanism in improving the robustness of AdaGC.
+TABLE IX: Test accuracies (%) Obtained With and Without Mixup
 Random Dominant